### v1

# Finding product sets in some classes of amenable groups

Abstract.
In [KMRR23b], using methods from ergodic theory, a longstanding conjecture of Erdős (see [Erd73, Page 305]) about sumsets in large subsets of the natural numbers was resolved. In this paper, we extend this result to several important classes of amenable groups, including all finitely generated virtually nilpotent groups, and all abelian groups $(G,+)$ with the property that the subgroup $2G:=\{g+g:g\in G\}$ has finite index. We prove that in any group $G$ from the above classes, any $A\subset G$ with positive upper Banach density contains a shifted product set of the form $\{tb_{i}b_{j}\colon i<j\}$, for some infinite sequence $(b_{n})_{n\in\mathbb{N}}$ and some $t\in G$. In fact, we show this result for all amenable groups that posses a property which we call square absolute continuity. Our results provide answers to several questions and conjectures posed in [KMRR23a].
1. Introduction
In [KMRR23b], Bryna Kra, Joel Moreira, Florian K. Richter and Donald Robertson, using methods from ergodic theory, proved that every subset $A$ of the positive integers with positive upper Banach density contains $\{b_{1}+b_{2}+t:b_{1}\neq b_{2}\in B\}$ for some infinite set $B\subset A$ and some $t\in\mathbb{N}$. This resolved a longstanding conjecture of Erdős (see [Erd73, Page 305]).
A natural question to ask is whether this result generalizes to other countable groups, such as $\mathbb{Z}^{d}$ for $d\geq 2$ or the discrete Heisenberg group, for example. The purpose of this paper is to extend the result in [KMRR23b] to several important classes of amenable groups, including all finitely generated virtually nilpotent groups and all abelian groups $(G,+)$ with the property that the subgroup consisting of the elements $2g:=g+g$, where $g\in G$, has finite index. To this end, we first extend the result to all amenable groups satisfying a property that we call square absolute continuity (see 1.4). Then we show that our result applies to the aforementioned classes of groups, by showing that they are (virtually) square absolutely continuous. Our main results provide partial answers to some questions and conjectures posed in [KMRR23a] regarding product sets in large subsets of amenable groups.
Throughout, let $G$ denote a countable group. Let us start with some basic definitions.
Definition 1.1.
Let $(G,\cdot)$ be a group. A sequence $\Phi=(\Phi_{N})_{N\in\mathbb{N}}$ of finite subsets of $G$ is:
  • a left Følner sequence, if it satisfies
    $\lim_{N\to\infty}\frac{|g\Phi_{N}\cap\Phi_{N}|}{|\Phi_{N}|}=1\text{\quad or equivalently, \quad}\lim_{N\to\infty}\frac{|g\Phi_{N}\triangle\Phi_{N}|}{|\Phi_{N}|}=0$
    for any $g\in G$, and
  • a right Følner sequence, if it satisfies
    $\lim_{N\to\infty}\frac{|\Phi_{N}g\cap\Phi_{N}|}{|\Phi_{N}|}=1\text{\quad or equivalently, \quad}\lim_{N\to\infty}\frac{|\Phi_{N}g\triangle\Phi_{N}|}{|\Phi_{N}|}=0$
    for any $g\in G$.
If both conditions are satisfied, then $\Phi$ is a two-sided Følner sequence.
We remark that if a group admits a left (or right) Følner sequence, then it admits a two-sided Følner sequence. Amenable groups, which are the central object of our study, are defined as follows:
Definition 1.2.
A group $(G,\cdot)$ is called amenable if it admits a left Følner sequence.
The most common example of an amenable group is $\mathbb{Z}^{d}$, for any $d\in\mathbb{N}$. Other examples of amenable groups are finite groups, abelian groups, solvable groups and finitely generated groups of subexponential growth. In addition, products of amenable groups, and virtually amenable groups are amenable. Følner sequences are useful to define notions of density in amenable groups.
Definition 1.3.
Let $(G,\cdot)$ be an amenable group, $\Phi$ a left (right) Følner sequence, and let $A\subset G$. Then the left (right) upper density of $A$ with respect to $\Phi$ is defined as
$\overline{d}_{\Phi}(A):=\limsup_{N\to\infty}\frac{|A\cap\Phi_{N}|}{|\Phi_{N}|}.$
We say that $A$ has positive left (right) upper Banach density if it has positive left (right) upper density with respect to some left (right) Følner sequence. We also say that $A$ has positive upper Banach density if it has positive upper density with respect to some two-sided Følner sequence.
Note that if $G$ is an amenable group, and $A\subset G$ has positive left upper Banach density, then this does not necessarily mean that $A$ has positive right upper Banach density.
Given a group $(G,\cdot)$, for any $g,h\in G$, Given a sequence $B=(b_{n})_{n\in\mathbb{N}}\subset G$, we define
[…]
and
$B\odot B:=\{b_{i}b_{j}:i\neq j\}.$
[…]
Definition 1.4.
Let $G$ be an amenable group and $\phi:G\to G$ be a map. We say that $G$ is $\phi$-absolutely continuous if $G$ admits two Følner sequences $\Phi=(\Phi_{N})_{N\in\mathbb{N}}$ and $\Psi=(\Psi_{N})_{N\in\mathbb{N}}$ satisfying the following: for any $\varepsilon>0$ there exists some $\delta>0$ such that for any $u:G\to[0,1]$ satisfying
$\limsup_{N\to\infty}\frac{1}{|\Phi_{N}|}\sum_{g\in\Phi_{N}}u(g)<\delta,$
we have that
$\limsup_{N\to\infty}\frac{1}{|\Psi_{N}|}\sum_{g\in\Psi_{N}}u(\phi(g))<\varepsilon.$
If, in particular, $\phi=s_{G}$, then we say that $G$ is square absolutely continuous.
1.1. Main results
The first main theorem of this paper is the following:
Theorem 1.5.
Let $G$ be a square absolutely continuous group and $A\subset G$ with positive left upper Banach density. Then there exist an infinite sequence $B=(b_{n})_{n\in\mathbb{N}}\subset A$ and some $t\in G$ such that
[…]
Theorem 1.5 provides a positive answer to [KMRR23a, Question 5.16], with the extra assumption that $G$ is square absolutely continuous, and under the weaker assumption that the set $A$ has positive left upper Banach density, instead of positive upper Banach density.
Remark 1.6.
Note that Theorem 1.5 immediately implies an analogous result for right upper Banach density instead of left upper Banach density. Indeed, through the map $g\mapsto g^{-1}$, Theorem 1.5 is equivalent to the assertion that for any $A\subset G$ with positive right upper Banach density, there exists an infinite sequence $C=(c_{n})_{n\in\mathbb{N}}\subset A$ and some $r\in G$ such that
[…]
Before continuing, let us recall the following definitions for a group $G$:
  • $G$ is nilpotent if its lower central series is finite, that is to say
    $G=G_{0}\triangleright G_{1}\triangleright\dots\triangleright G_{n}=\{e_{G}\},$
    where $G_{i+1}:=[G_{i},G]$ is the commutator group of $G_{i}$ and $G$, i.e., the subgroup of $G$ generated by the elements of the form $hgh^{-1}g^{-1}$, where $h\in G_{i},g\in G$.
  • $G$ is finitely generated if there exist $g_{1},\dots,g_{n}\in G$ such that any element of $G$ can be written as product of $g_{1},\dots,g_{n}$.
  • $G$ is torsion-free if it does not have any non-trivial element of finite rank, that is to say, for any $g\in G$ with $g\neq e_{G}$ and any $n\in\mathbb{N}$, we have $g^{n}\neq e_{G}$.
  • If $P$ is a property of groups, then we say that a group $G$ is virtually $P$ if it has a finite-index subgroup that has the property $P$.
In any finitely generated nilpotent group $G$, there exist some $s\in\mathbb{N}$ (depending on the degree of nilpotency and the number of generators of $G$), some $a_{i}\in G$ and some functions $t_{i}:G\to\mathbb{Z}$, for $1\leq i\leq s$, such that any $x\in G$ can be written as $x=a_{1}^{t_{1}(x)}\cdots a_{s}^{t_{s}(x)}$. The $s$-tuple $(a_{1},\dots,a_{s})$ is a Mal’cev basis and the $s$-tuple $(t_{1},\dots,t_{s})$ is a Mal’cev coordinate system with respect to this Mal’cev basis. If $G$ is also torsion-free, then the coordinate maps are injective and hence we can identify $G$ with $\mathbb{Z}^{s}$ and it is convenient to also identify any $x\in G$ with its coordinates $(t_{1}(x),\dots,t_{s}(x))\in\mathbb{Z}^{s}$. The above facts about Mal’cev bases can be found in [KM79, Chapter 17.2].
Theorem 1.7.
Every torsion-free finitely generated nilpotent group is square absolutely continuous.
Combining Theorems 1.5 and 1.7 we have that every torsion-free finitely generated nilpotent group satisfies the conclusion of Theorem 1.5. In fact, we prove the following slight strengthening:
Corollary 1.8.
Let $G$ be a torsion-free finitely generated nilpotent group and $A\subset G$ with positive left upper Banach density. Then there exist an infinite sequence $B=(b_{n})_{n\in\mathbb{N}}\subset A$ and some $t\in G$ such that
[…]
Moreover, given a Mal’cev coordinate system $(t_{1},\dots,t_{s})$ on $G$, we can choose $B$ so that the following holds: for any finite set $C\subset\mathbb{Z}$ and any $1\leq i\leq s$, the set $\{b\in B\colon t_{i}(b)\in C\}$ is finite.
Furthermore, we are able to extend the first statement of Corollary 1.8 to all finitely generated virtually nilpotent groups.
Corollary 1.9.
Let $G$ be a finitely generated virtually nilpotent group and $A\subset G$ with positive left upper Banach density. Then there exist $g\in G$, an infinite sequence $B=(b_{n})_{n\in\mathbb{N}}\subset g^{-1}A$, and some $t\in G$ such that
[…]
In particular, this holds for the group $UT(n,F)$111 $UT(n,F)$ is the unitriangular $n\times n$ matrix group with entries from $F$. Note that $U(3,\mathbb{Z})$ is the well-known $3\times 3$ Heisenberg group., where $n\in\mathbb{N}$ and $F$ is any infinite, finitely generated field.
Remark 1.10.
Note that the class of finitely generated virtually nilpotent groups coincides, in view of Gromov’s theorem, with the class of finitely generated groups of polynomial growth.
The final part of this subsection is concerned with sumsets in abelian groups. Let $G=(G,+)$ be an abelian group. We write $2g$ to denote the element $g+g$, for any $g\in G$. Moreover, we refer to the map $s_{G}:G\to G$ as the doubling map and its image is now the subgroup of $G$ consisting of all elements of the form $2g$, where $g\in G$. We denote this subgroup by $2G$, i.e., $s_{G}(G)=2G$, and we often refer to it as the doubling subgroup.
In [KMRR23a, Conjecture 5.14], it is conjectured that in any countable abelian group $G$, every set of positive upper Banach density contains a set of the form $B\oplus B+t=\{b_{1}+b_{2}+t:b_{1}\neq b_{2}\in B\}$ for an infinite set $B\subset G$ and some $t\in G$. It follows from Corollary 1.9 that this conjecture holds under the additional assumption that $G$ is finitely generated, and moreover, it extends [KMRR23b, Theorem 1.2] from $(\mathbb{N},+)$ to all finitely generated abelian groups. In fact, we are able to verify [KMRR23a, Conjecture 5.14], and hence extend [KMRR23b, Theorem 1.2] to an even larger collection of abelian groups, that contains all the finitely generated abelian groups along with some infinitely generated ones. The following theorem allows us to do so:
Theorem 1.11.
Every abelian group whose doubling subgroup has finite index is square absolutely continuous.
The following corollary is an obvious consequence of Theorems 1.5 and 1.11:
Corollary 1.12.
Let $(G,+)$ be an abelian group such that $2G$ is a finite-index subgroup of $G$, and let $A\subset G$ with positive upper Banach density. Then there exist an infinite set $B\subset A$ and some $t\in G$ such that
$B\oplus B\subset A-t.$
In particular, this holds for:
  • all finitely generated abelian groups,222 It is not hard to check that in finitely generated abelian groups, the doubling subgroup has finite-index. and
[…]
In the semigroup $(\mathbb{N},\times)$ the sub-semigroup $\{n^{2}\colon n\in\mathbb{N}\}$ has infinite index (and zero density), and unfortunately, our methods do not allow us to handle this case.
We remark that Corollary 1.12 can also be proved independently of Theorem 1.5, meaning that, by slightly modifying the proof of Theorem 1.5, one can directly obtain the result for abelian groups with finite-index doubling subgroup without showing that such groups are square absolutely continuous.
1.2. More product sets and open questions
The following remark shows that in the formulation of Theorem 1.5 one can replace left shifts with right shifts and the statement remains true.
Remark 1.13.
Let $G$ and $A$ be as in Theorem 1.5. Then there exist some $t\in G$ and some $B=(b_{n})_{n\in\mathbb{N}}\subset tAt^{-1}$ such that
[…]
To see why, let $B^{\prime}\subset A$ and $t\in G$ such that $B^{\prime}\hskip 2.13394pt\mathrlap{\mkern 3.3mu{{\cdot}}}{\triangleleft}\hskip 2.84526ptB^{\prime}\subset t^{-1}A$, as guaranteed by Theorem 1.5, and then let $B=tB^{\prime}t^{-1}$.
[…]
Product sets in sets of positive left upper Banach density
[…]
True (Theorem 1.5; for $\overline{d}_{\Phi}(A)>0$ for some left Følner $\Phi$)
[…]
True (Remark 1.13; for $\overline{d}_{\Phi}(A)>0$ for some left Følner $\Phi$)
[…]
False (Example 6.1; with $\overline{d}_{\Phi}(A)>0$ for some left Følner $\Phi$)
[…]
False (Example 6.2; with $\overline{d}_{\Phi}(A)=1$ for some left Følner $\Phi$)
[…]
False (Example 6.3; with $\overline{d}_{\Phi}(A)=1$ for some left Følner $\Phi$)
[…]
It remains interesting to ask whether product sets of the form $B\odot B$ can be found in sets with positive upper Banach density. Unfortunately, our methods here are insufficient to handle this case. In this spirit, we conclude this section with the two questions below. We remark that the second one is a special case of [KMRR23a, Question 5.17].
Question 1.14.
Let $G$ be a square absolutely continuous group and $A\subset G$ be a set of positive upper Banach density. Is it true that there exists some set $B\subset G$ such that
$B\odot B\subset t^{-1}A\cup Ar^{-1}$
for some $t,r\in G$?
Question 1.15.
Let $G$ and $A$ be as in 1.14. Is it true that there exists some set $B\subset G$ such that
$B\odot B\subset t^{-1}Ar^{-1}$
for some $t,r\in G$?
1.3. Proof ideas
To prove Theorem 1.5, we follow an ergodic-theoretic approach and we employ ideas similar to the ones used in [KMRR23b] in the setting of $(\mathbb{N},+)$. This approach is based on methods that were introduced in [KMRR22] to generalize another sumset conjecture of Erdős, which was initially proved in [MRR19] by Moreira, Richter and Robertson. However, the generality of the setting of amenable groups compared to $(\mathbb{N},+)$ causes several issues and complications that we need to handle differently. These issues, along with the new ideas we develop to deal with them, are briefly discussed below.
After translating Theorem 1.5 into a dynamical statement (see Theorem 3.8), we reduce the problem to finding certain dynamical configurations given by limit points of orbits of ergodic measure-preserving $G$-actions, called Erdős progressions (see 3.2). The natural environment in which one can study such progressions is the Kronecker factor of a system, as Erdős progressions are simply $3$-term arithmetic progressions there.
One of the main obstructions we had to overcome in our proof is the lack of commutativity of $G$. The most notable among the issues that this leads to is that the Kronecker factor does not have the structure of an abelian group, but instead, it is a homogeneous space $Z=K/H$, for some compact group $K$. This makes the study of Erdős progressions more technically challenging. To be more precise, the abelian nature of the Kronecker factor in the setting of $(\mathbb{N},+)$-actions is heavily used in [KMRR23b]. Consequently, due to the absence of commutativity in our case, many of the techniques in [KMRR23b] do not generalize easily to our setting. Another difficulty that arises in non-commutative groups is the erratic behavior of the set of squares $G^{2}$. In particular, orbits of points along $G^{2}$ may be trapped in zero-measure regions, which causes serious trouble in finding Erdős progressions. The assumption that $G$ is square absolutely continuous is critical in avoiding this scenario. In addition, we need an extension of a result of Host and Kra ([HK09, Proposition 6.1]) concerning actions of $(\mathbb{N},+)$, to the more general setting of amenable group actions (see Lemma 3.5, proof in Appendix A).
Acknowledgements. We would like to thank Joel Moreira and Florian K. Richter for their insightful suggestions, beneficial comments and constant support throughout the writing of this paper. We also want to thank Ethan Ackelsberg and Felipe Hernández Castro for their useful comments on earlier drafts of this paper.
The first author gratefully acknowledges support from the Swiss National Science Foundation grant TMSGI2-211214. The second author is supported by the Warwick Mathematics Institute Centre for Doctoral Training, and gratefully acknowledges funding by University of Warwick’s Chancellors’ International Scholarship scheme.
2. Preliminaries
In this section, we state all the preliminaries that will be useful in the rest of the paper regarding classic notions and theorems of ergodic theory of actions of amenable groups. So, for the rest of the section, $G$ denotes an arbitrary countable and discrete amenable group.
Basics on $G$-systems: Given a compact metric space $X=(X,d_{X})$, a continuous action $T=(T_{g})_{g\in G}$ of $G$ on X is a collection of continuous functions $T_{g}:X\to X$ such that for any $g_{1},g_{2}\in G$, $T_{g_{1}}\circ T_{g_{2}}=T_{g_{1}g_{2}}$. Given such an action, we call the pair $(X,T)$ a topological $G$-system.
Given a topological $G$-system $(X,T)$ and a point $x\in X$, we define its orbit as $\mathcal{O}_{T}(x)=\{T_{g}x:g\in G\},$ and we say that the point is transitive if $\mathcal{O}_{T}(x)$ is dense in $X$.
Fix a topological $G$-system $(X,T)$. Let $M(X)$ denote the space of Borel probability measures on $X$, equipped with the ${}^{\ast}$ topology, which is compact and metrizable. A measure $\mu\in M(X)$ is said to be $T$-invariant, if it is invariant under $T_{g}$ for all $g\in G$. The subset of $M(X)$ consisting of $T$-invariant measures is denoted by $M^{T}(X)$, and it is a closed and convex subset of $M(X)$. The Borel $\sigma$-algebra on $X$ is denoted by $\mathscr{B}_{X}$ or just $\mathscr{B}$, if no confusion may arise.
For $\mu\in M^{T}(X)$, the action $T$ on the Borel probability space $(X,\mu)$ is called a measure-preserving $G$-action and $(X,\mu,T)$ is called a measure-preserving $G$-system. Note that we omit writing the symbol for the $\sigma$-algebra, and from now on, whenever this happens, the implied $\sigma$-algebra will be the Borel. For simplicity, we refer to the above as $G$-actions, and $G$-systems, respectively.
Given a $G$-system $(X,\mu,T)$, one can define an action, which by abuse of notation will again be denoted by $T=(T_{g})_{g\in G}$, of $G$ on $L^{2}(X)$ by $T_{g}:L^{2}(X)\to L^{2}(X)$, $T_{g}f=f\circ T_{g}$. It is not hard see that for all $g\in G$, $T_{g}$ is an isometry of $L^{2}(X)$. Note also that since $G$ acts from the left on $X$, then $G$ acts from the right on $L^{2}(X)$.
We remark that we are only considering $G$-systems where $G$ acts on the prescribed space from the left, and then any associated Følner will be considered left, without mentioning it, unless it is necessary.
Product $G$-system: Given two $G$-systems $(X,\mu,T)$ and $(Y,\nu,S)$, we define the product $G$-system $(X\times Y,\mu\times\nu,T\times S)$, where the underlying $\sigma$-algebra is the product of the Borel $\sigma$-algebras on $X$ and $Y$, which coincides with the Borel $\sigma$-algebra on $X\times Y$, and the action is $T\times S=(T_{g}\times S_{g})_{g\in G}$.
Factors of $G$-systems: Given two $G$-systems $(X,\mu,T)$ and $(Y,\nu,S)$, we say that $(Y,\nu,S)$ is a factor of $(X,\mu,T)$ if there exists a measurable map $\pi:X\to Y$, which we call a factor map, satisfying: $\mu(\pi^{-1}E)=\nu(E)$ for any measurable $E\subset Y$ and for any $g\in G$, $\pi\circ T_{g}=S_{g}\circ\pi$ $\mu$-almost everywhere on $X$. When the former is true, we say that $\nu$ is the push-forward of $\mu$ under $\pi$, and we write $\pi\mu=\nu$. When, additionally, the factor map $\pi$ is continuous and $\pi\circ T_{g}=S_{g}\circ\pi$ holds everywhere on $X$ for any $g\in G$, we say that $\pi$ is a continuous factor map and $(Y,\nu,S)$ is a continuous factor of $(X,\mu,T)$.
Ergodicity and ergodic theorems for $G$-systems: A $G$-system $(X,\mu,T)$ is called ergodic if for any measurable set $A$ the following holds:
$T_{g}^{-1}A=A\text{ for all }g\in G\implies\mu(A)=0\text{ or }\mu(A)=1.$
Given a $G$-system $(X,\mu,T)$, let $\mathscr{A}$ be a sub-$\sigma$-algebra of $\mathscr{B}$. For $f\in L^{2}(X,\mu)$, the conditional expectation of $f$ on $\mathscr{A}$, denoted by $\mathbb{E}_{\mu}(f\>|\>\mathscr{A})$, is defined as the orthogonal projection of $f$ on the closed subspace $L^{2}(X,\mathscr{A},\mu)$ of $L^{2}(X,\mu)$. We also denote the sub-$\sigma$-algebra of the $T$-invariant sets by $\mathcal{I}=\mathcal{I}(T)$; that is,
$\mathcal{I}=\mathcal{I}(T):=\{E\in\mathscr{B}:T_{g}^{-1}E=E\text{ for all }g\in G\}.$
Theorem 2.1 (Mean Ergodic Theorem for $G$-systems, see [Gla03, Theorem 3.33]).
Let $(X,\mu,T)$ be a $G$-system, and let $\Phi$ be a Følner sequence. Then, for any $f\in L^{2}(X)$,
$\frac{1}{|\Phi_{N}|}\sum_{g\in\Phi_{N}}T_{g}f\to\mathbb{E}_{\mu}(f\>|\>\mathcal{I})$
as $N\to\infty$ in $L^{2}(X)$. In addition, if the system is ergodic, the ergodic averages above converge to $\int_{X}f~{}\mathrm{d}\mu$.
Measure disintegration and ergodic decomposition: When $X$ is a compact metric space, the space $M(X)$ of Borel probability measures on $X$ can be endowed with a $\sigma$-algebra $\mathcal{M}$ such that the space $(M(X),\mathcal{M})$ is a standard Borel space. The following theorem about disintegrations of measures is very useful.
Theorem 2.2 (Disintegration of measures, see [HK18, Chapter 2, Section 2.5]).
Let $X$ be a compact metric space, $\mathscr{B}$ the Borel $\sigma$-algebra on $X$ and $\mu$ a probability measure on $(X,\mathscr{B})$. Let also $\mathscr{D}$ be a sub-$\sigma$-algebra of $\mathscr{B}$. Then there is a $(\mathscr{D},\mu)$-almost everywhere defined and measurable map $(X,\mathscr{D})\to(M(X),\mathcal{M})$, $x\mapsto\mu_{x}$ with the following properties:
  • For every $f\in L^{1}(X,\mu)$, the function $x\mapsto\int_{X}f~{}\mathrm{d}\mu_{x}$ is in $L^{1}(X,\mathscr{D},\mu)$, and for all $D\in\mathscr{D}$, we have $\int_{D}f~{}\mathrm{d}\mu=\int_{D}\big{(}\int_{X}f~{}\mathrm{d}\mu_{x}\big{)}~{}\mathrm{d}\mu(x).$ In particular, this implies that $\int_{X}f~{}\mathrm{d}\mu_{x}=\mathbb{E}_{\mu}(f\>|\>\mathscr{D})(x)$ for $(\mathscr{D},\mu)$-almost every $x\in X$.
  • $\mu_{x}([x]_{\mathscr{D}})=1$, where $[x]_{\mathscr{D}}=\cap_{x\in D\in\mathscr{D}}D.$
The map satisfying the above properties is unique modulo $(\mathscr{D},\mu)$-null sets, and is called the disintegration of the measure $\mu$ over the sub-$\sigma$-algebra $\mathscr{D}$. In that case, we write $\mu=\int_{X}\mu_{x}~{}\mathrm{d}\mu(x)$.
In this paper, we will extensively make use of disintegrations over (continuous) factor maps. Let $\pi:(X,\mu,T)\to(Y,\nu,S)$ be a factor map between two $G$-systems. In the setting of Theorem 2.2, let $\mathscr{D}=\pi^{-1}(\mathscr{B}_{Y})$, where $\mathscr{B}_{Y}$ is the Borel $\sigma$-algebra on $Y$. Then Theorem 2.2 gives a disintegration $y\mapsto\mu_{y}$ defined on $Y$, which is unique up to $\nu$-null measure sets, and satisfies the following:
  • for $\nu$-almost every $y\in Y$,
    (2.1) $\mathbb{E}_{\mu}(f\>|\>Y)(y):=\mathbb{E}_{\mu}(f\>|\>\mathscr{D})(y)=\int_{Y}f~{}\mathrm{d}\mu_{y},$
  • for $\nu$-almost every $y\in Y$, $\mu_{y}(\pi^{-1}(\{y\}))=1$, and finally,
  • for any $g\in G$ and for $\nu$-almost every $y\in Y$, $(T_{g})\mu_{y}=\mu_{S_{g}y}$.
Let $(X,\mu,T)$ be a $G$-system. Consider the (unique) disintegration of $\mu$ with respect to $\mathcal{I}=\mathcal{I}(T)$ given by Theorem 2.2. This disintegration is called the ergodic decomposition of $\mu$. Equivalently, we say that the disintegration $x\mapsto\mu_{x}$ is the ergodic decomposition of $\mu$, if for any $f:X\to\mathbb{C}$ measurable and bounded,
(2.2) $\int_{X}f~{}\mathrm{d}\mu_{x}=\mathbb{E}_{\mu}(f\>|\>\mathcal{I})(x)$
holds for $(\mathcal{I},\mu)$-almost every $x\in X$.
Theorem 2.3 (Ergodic decomposition of $G$-systems, see [Gla03, Theorem 3.22]).
If $(X,\mu,T)$ is a $G$-system as above, then for $(\mathcal{I},\mu)$-almost every $x\in X$, the measure $\mu_{x}$ is $T$-invariant and the system $(X,\mu_{x},T)$ is ergodic.
Generic points and the support of a measure: In addition, we will need the notion of generic points:
Definition 2.4.
Let $(X,\mu,T)$ be a $G$-system and let $\Phi$ be a Følner sequence. A point $a\in X$ is called generic for $\mu$ along $\Phi$ if for all $f\in C(X)$ we have
$\lim_{N\to\infty}\frac{1}{|\Phi_{N}|}\sum_{g\in\Phi_{N}}f(T_{g}a)=\int_{X}f~{}\mathrm{d}\mu$
or equivalently if
$\lim_{N\to\infty}\frac{1}{|\Phi_{N}|}\sum_{g\in\Phi_{N}}\delta_{T_{g}a}=\mu$
where $\delta_{x}$ is the Dirac mass at $x\in X$ and the limit is in the ${}^{\ast}$ topology. If $a$ is generic for $\mu$ along $\Phi$, then we denote this by $a\in\texttt{{gen}}(\mu,\Phi)$.
Moreover, we will need the notion of the support of a measure. The support of a Borel probability measure $\mu$ on a compact metric space $X$ is the smallest closed full-measure subset of $X$ and is denoted by $\texttt{{supp}}(\mu)$. We will need the following lemma, which says that generic points for a measure have dense orbit in the support of the measure. Its proof is quite standard, and we only include it for completeness.
Lemma 2.5.
Let $(Y,\nu,S)$ be a $G$-system and let $y,w\in Y$. If $y\in\texttt{{gen}}(\nu,\Phi)$ for some Følner sequence $\Phi$, and $w\in\texttt{{supp}}(\nu)$, then $S_{g_{n}}y\to w$, for some infinite sequence $(g_{n})_{n\in\mathbb{N}}$ in $G$.
Proof.
Fix a compatible metric on $X$ and let $\texttt{{B}}(w,\varepsilon)$ be the open ball centered at $w$ with radius $\varepsilon>0$ with respect to this metric. By Urysohn’s lemma, for every $\varepsilon>0$ there exists a continuous function $f:X\to[0,1]$ with $f=1$ on $\texttt{{B}}(w,\varepsilon/2)$ and $f=0$ outside $\texttt{{B}}(w,\varepsilon)$. Since $w\in\texttt{{supp}}(\nu)$, it follows that $\int_{Y}f~{}\mathrm{d}\nu>0$. Now, using that $y\in\texttt{{gen}}(\nu,\Phi)$, we have that
$\lim_{N\to\infty}\frac{1}{|\Phi_{N}|}\sum_{g\in\Phi_{N}}f(S_{g}y)=\int_{X}f~{}\mathrm{d}\nu>0,$
which implies that $S_{g}y\in\texttt{{B}}(w,\varepsilon)$ for infinitely many $g\in G$. The result then follows. ∎
We will also make use of the following result of Lindenstrauss:
Proposition 2.6 (see [Lin99, Theorem 1.2 and Proposition 1.4]).
Let $(X,\mu,T)$ be a system and $\Phi$ be a Følner sequence in $G$. Then there is a subsequence $\Psi$ of $\Phi$ such that for all $f\in L^{1}(\mu)$,
$\lim_{N\to\infty}\frac{1}{|\Psi_{N}|}\sum_{g\in\Psi_{N}}T_{g}f(x)=\mathbb{E}_{\mu}(f\>|\>\mathcal{I})(x)$
for $\mu$-almost every $x\in X$.
The next two lemmas follow easily from Proposition 2.6:
Lemma 2.7.
Let $(X,\mu,T)$ be an ergodic $G$-system. Then for any Følner sequence $\Phi$ there exists some subsequence $\Psi$ such that $\mu$-almost every $x\in X$ is in $\texttt{{gen}}(\mu,\Psi)$.
Lemma 2.8.
Let $(X,\mu,T)$ be a $G$-system, let $\Phi$ be a Følner sequence, and let $x\mapsto\mu_{x}$ be the ergodic decomposition of $\mu$. Then there exists some subsequence $\Psi$ of $\Phi$ such that $\mu$-almost every $x\in X$ is in $\texttt{{gen}}(\mu_{x},\Psi)$.
Finally, it will be useful to have the following generalisation of [Fur81, Proposition 3.9], whose proof is again the same as for actions of $(\mathbb{N},+)$, but we include it for the convenience of the reader.
Lemma 2.9.
Let $G$ be an amenable group, let $(X,\mu,T)$ be an ergodic $G$-system, and let $a\in X$ be a point such that $\mu$ is supported on $\overline{\mathcal{O}_{T}(a)}$. Then there exists some Følner sequence $\Psi$ such that $a\in\texttt{{gen}}(\mu,\Psi)$.
Proof.
By Lemma 2.7, there exists some $x_{0}\in\overline{\mathcal{O}_{T}(a)}$ that is generic for $\mu$ along some Følner sequence $\Phi$. Let $\mathcal{F}=(f_{k})_{k\in\mathbb{N}}$ be a dense subset of $(C(X),\|\cdot\|_{\infty})$ and let $(\Phi_{N_{n}})_{n\in\mathbb{N}}$ be a subsequence of $\Phi$ such that for every $n\in\mathbb{N}$ and for every $j=1,2,\dots,n$,
$\bigg{|}\frac{1}{|\Phi_{N_{n}}|}\sum_{g\in\Phi_{N_{n}}}f_{j}(T_{g}x_{0})-\int_{X}f_{j}~{}\mathrm{d}\mu\bigg{|}<\frac{1}{n}.$
Since $x_{0}\in\overline{\mathcal{O}_{T}(a)}$, there exists some $(g_{n})_{n\in\mathbb{N}}\subset G$ such that $T_{g_{n}}a\to x_{0}$, so that we may assume that the equation above holds if we substitute $x_{0}$ with $T_{g_{n}}a$. Consider the Følner sequence $\Psi=(\Psi_{n})$ given by $\Psi_{n}=\Phi_{N_{n}}g_{n}$. It follows that for every $n\in\mathbb{N}$ and any $j=1,2,\dots,n$,
$\bigg{|}\frac{1}{|\Psi_{n}|}\sum_{g\in\Psi_{n}}f_{j}(T_{g}a)-\int_{X}f_{j}~{}\mathrm{d}\mu\bigg{|}<\frac{1}{n}.$
Since $\mathcal{F}$ is dense in $C(X)$ the conclusion follows as before. ∎
Kronecker factor and the Jacobs-de Leeuw-Glicksberg decomposition: Let $(X,\mu,T)$ be a $G$-system. A function $f\in L^{2}(X)$ is called:
  • compact, if $\overline{\{T_{g}f:g\in G\}}$ is compact with respect to the strong topology on $L^{2}(X)$.
  • weak-mixing, if for any Følner sequence $\Phi$, and any $f^{\prime}\in L^{2}(X)$,
    $\lim_{N\to\infty}\frac{1}{|\Phi_{N}|}\sum_{g\in\Phi_{N}}|\langle T_{g}f,f^{\prime}\rangle|=0.$
[…]
In case that $G$ is an amenable group, the Jacobs-de Leeuw-Glicksberg decomposition theorem applies, stating that these two components give a decomposition of $L^{2}(X)$.
Theorem 2.10 (Jacobs-de Leeuw-Glicksberg decomposition, see [KL16, Theorem 2.24]).
If $(X,\mu,T)$ is a $G$-system, then
$L^{2}(X)=\mathcal{H}_{\textup{c}}\oplus\mathcal{H}_{\textup{wm}}.$
Now we will give a description of the factor of $(X,\mu,T)$ corresponding to the subspace $\mathcal{H}_{\textup{c}}$ of $L^{2}(X)$. Note that if $f\in\mathcal{H}_{\textup{c}}$, then for all $g\in G$, $T_{g}f\in\mathcal{H}_{\textup{c}}$, so $\mathcal{H}_{\textup{c}}$ in invariant under the action of $T$ on $L^{2}(X)$. Let $\mathcal{A}$ be the smallest $\sigma$-algebra with respect to which all functions in $\mathcal{H}_{\textup{c}}$ are measurable. Then $\mathcal{A}$ is $T$-invariant $\sigma$-algebra contained in the Borel $\sigma$-algebra of $X$. Therefore, the system $(X,\mathcal{A},\mu,T)$ is a factor of the original system, with the factor map being the identity $\textup{id}:X\to X$. This factor is called the Kronecker factor of $(X,\mu,T)$.
Our goal now is to give a nice algebraic description of the Kronecker factor when the $G$-system $(X,\mu,T)$ is ergodic.
Proposition 2.11.
[Mac64, Theorem 1] Let $(X,\mu,T)$ be an ergodic $G$-system. Then there exist a compact group $K$, a closed subgroup $H$ of $K$ and a continuous group homomorphism $\alpha:G\to K$ with dense image such that the Kronecker factor of $(X,\mu,T)$ is measurably isomorphic to the $G$-system $(Z,m,R)$ given as follows: $Z=K/H$ is a homogeneous space, $m$ is the normalized Haar measure on $Z$ and $R=(R_{g})_{g\in G}$, where for each $g\in G$, $R_{g}:Z\to Z$ is given by $R_{g}(z)=\alpha(g)z$. The system $(Z,m,R)$ is called a rotation on the homogeneous space $Z$.
This proposition allows us to identify the Kronecker factor with a rotation on a homogeneous space, whenever $(X,\mu,T)$ is ergodic.
Characteristic factors for $G$-systems: The notion of characteristic factors will play a fundamental role later in one of our proofs. Characteristic factors are relevant to the ergodic averages in question, in each different problem. Here we have the following theorem for the characteristic factors with respect to some double averages that will concern us.
Theorem 2.12.
Let $(X,\mu,T)$ be an ergodic $G$-system, let $(Z,m,R)$ be its Kronecker factor and $\Phi$ be a Følner sequence. Then for any $f_{1},f_{2}\in L^{\infty}(X)$, we have
$\lim_{N\to\infty}\frac{1}{|\Phi_{N}|}\sum_{g\in\Phi_{N}}T_{g}f_{1}\otimes T_{g}f_{2}=\lim_{N\to\infty}\frac{1}{|\Phi_{N}|}\sum_{g\in\Phi_{N}}T_{g}\mathbb{E}_{\mu}(f_{1}\>|\>Z)\otimes T_{g}\mathbb{E}_{\mu}(f_{2}\>|\>Z)$
in $L^{2}(X\times X,\mu\times\mu)$. We say then that the characteristic factor for the averages in the left-hand side of the above is the Kronecker.
The proof of this theorem will follow easily from the following lemma.
Lemma 2.13.
Let $(X,\mu,T)$ be a $G$-system. Then
$\mathcal{H}_{\textup{wm}}(T)\otimes L^{2}(X)\subset\mathcal{H}_{\textup{wm}}(T\times T)\>\text{ and }\>L^{2}(X)\otimes\mathcal{H}_{\textup{wm}}(T)\subset\mathcal{H}_{\textup{wm}}(T\times T).$
Proof.
We will only prove the first inclusion, as the second follows in an analogous way. Let $\Phi$ be any Følner sequence, and let $f_{1}\in\mathcal{H}_{\textup{wm}}(T)$ and $f_{2}\in L^{2}(X)$. We may assume that $\|f_{1}\|_{2},\|f_{2}\|_{2}\leq 1$. We want to show that
(2.3) $\lim_{N\to\infty}\frac{1}{|\Phi_{N}|}\sum_{g\in\Phi_{N}}\Big{|}\langle(T_{g}\times T_{g})(f_{1}\otimes f_{2}),F\rangle_{L^{2}(\mu\times\mu)}\Big{|}^{2}=0$
for any $F\in L^{2}(X\times X,\mu\times\mu)$.
Let $F\in L^{2}(X\times X,\mu\times\mu)$ and $\varepsilon>0$. We may assume that $\|F\|_{L^{2}(\mu\times\mu)}=1$. Now, since finite linear combinations of functions of the form $f^{\prime}_{1}\otimes f^{\prime}_{2}$, where $f^{\prime}_{1},f^{\prime}_{2}\in L^{2}(X)$, form a dense subset of $L^{2}(X\times X,\mu\times\mu)$, we can find $F^{\prime}=\sum_{i=1}^{k}c_{i}(f_{1,i}^{\prime}\otimes f_{2,i}^{\prime})$ with $\|F^{\prime}\|_{L^{2}(\mu\times\mu)}\leq 1$ such that $\|F-F^{\prime}\|_{L^{2}(\mu\times\mu)}<\varepsilon/2$. Then by the Cauchy-Schwarz inequality we have
$\big{|}\langle(T_{g}\times T_{g})(f_{1}\otimes f_{2}),F\rangle_{L^{2}(\mu\times\mu)}\big{|}^{2}=\langle(T_{g}\times T_{g})(f_{1}\otimes f_{2}),F\rangle_{L^{2}(\mu\times\mu)}\langle F,(T_{g}\times T_{g})(f_{1}\otimes f_{2})\rangle_{L^{2}(\mu\times\mu)}$
$=\langle(T_{g}\times T_{g})(f_{1}\otimes f_{2}),F^{\prime}\rangle_{L^{2}(\mu\times\mu)}\langle F^{\prime},(T_{g}\times T_{g})(f_{1}\otimes f_{2})\rangle_{L^{2}(\mu\times\mu)}$
$\hskip 11.38092pt+\langle(T_{g}\times T_{g})(f_{1}\otimes f_{2}),F^{\prime}\rangle_{L^{2}(\mu\times\mu)}\langle F-F^{\prime},(T_{g}\times T_{g})(f_{1}\otimes f_{2})\rangle_{L^{2}(\mu\times\mu)}$
$\hskip 11.38092pt+\langle(T_{g}\times T_{g})(f_{1}\otimes f_{2}),F-F^{\prime}\rangle_{L^{2}(\mu\times\mu)}\langle F,(T_{g}\times T_{g})(f_{1}\otimes f_{2})\rangle_{L^{2}(\mu\times\mu)}$
$<\sum_{1\leq i,j\leq k}\overline{c_{i}}c_{j}\langle(T_{g}\times T_{g})(f_{1}\otimes f_{2}),f_{1,i}^{\prime}\otimes f_{2,i}^{\prime}\rangle_{L^{2}(\mu\times\mu)}\langle f_{1,j}^{\prime}\otimes f_{2,j}^{\prime},(T_{g}\times T_{g})(f_{1}\otimes f_{2})\rangle_{L^{2}(\mu\times\mu)}+\varepsilon$
$\leq\sum_{1\leq i,j\leq k}\overline{c_{i}}c_{j}\|f_{2,i}^{\prime}\|_{2}\|f_{1,j}^{\prime}\|_{2}\|f_{2,j}^{\prime}\|_{2}|\langle T_{g}f_{1},f_{1,i}^{\prime}\rangle|+\varepsilon.$
Therefore, using that $f_{1}$ is a weak-mixing function, we have that
$\limsup_{N\to\infty}\frac{1}{|\Phi_{N}|}\sum_{g\in\Phi_{N}}\big{|}\langle(T_{g}\times T_{g})(f_{1}\otimes f_{2}),F\rangle_{L^{2}(\mu\times\mu)}\big{|}^{2}=$
$=\sum_{1\leq i,j\leq k}\overline{c_{i}}c_{j}\|f_{2,i}^{\prime}\|_{2}\|f_{1,j}^{\prime}\|_{2}\|f_{2,j}^{\prime}\|_{2}\limsup_{N\to\infty}\frac{1}{|\Phi_{N}|}\sum_{g\in\Phi_{N}}|\langle T_{g}f_{1},f_{1,i}^{\prime}\rangle|+\varepsilon=\varepsilon.$
Since $\varepsilon>0$ was arbitrary, then (2.3) follows. The proof is complete. ∎
Proof of Theorem 2.12.
Let $\Phi$ be Følner sequence, and let $f_{1},f_{2}\in L^{2}(X)$. Then we write
$\frac{1}{|\Phi_{N}|}\sum_{g\in\Phi_{N}}T_{g}f_{1}\otimes T_{g}f_{2}$ $=\frac{1}{|\Phi_{N}|}\sum_{g\in\Phi_{N}}T_{g}(f_{1}-\mathbb{E}_{\mu}(f_{1}\>|\>Z))\otimes T_{g}(f_{2}-\mathbb{E}_{\mu}(f_{2}\>|\>Z))$
$+\frac{1}{|\Phi_{N}|}\sum_{g\in\Phi_{N}}T_{g}(f_{1}-\mathbb{E}_{\mu}(f_{1}\>|\>Z))\otimes T_{g}\mathbb{E}_{\mu}(f_{2}\>|\>Z)$
$+\frac{1}{|\Phi_{N}|}\sum_{g\in\Phi_{N}}T_{g}\mathbb{E}_{\mu}(f_{1}\>|\>Z)\otimes T_{g}(f_{2}-\mathbb{E}_{\mu}(f_{2}\>|\>Z))$
(2.4) $+\frac{1}{|\Phi_{N}|}\sum_{g\in\Phi_{N}}T_{g}\mathbb{E}_{\mu}(f_{1}|Z)\otimes T_{g}\mathbb{E}_{\mu}(f_{2}\>|\>Z).$
Note that the limits of all the terms above exist by the mean ergodic theorem (Theorem 2.1) applied to $T\times T$. By Theorem 2.10, the functions $f_{1}-\mathbb{E}_{\mu}(f_{1}\>|\>Z),f_{2}-\mathbb{E}_{\mu}(f_{2}\>|\>Z)$ are both weak-mixing. Then, by Lemma 2.13, the functions $(f_{1}-\mathbb{E}_{\mu}(f_{1}\>|\>Z))\otimes(f_{2}-\mathbb{E}_{\mu}(f_{2}\>|\>Z)),(f_{1}-\mathbb{E}_{\mu}(f_{1}\>|\>Z))\otimes\mathbb{E}_{\mu}(f_{2}\>|\>Z)$ and $\mathbb{E}_{\mu}(f_{1}\>|\>Z)\otimes(f_{2}-\mathbb{E}_{\mu}(f_{2}\>|\>Z))$ are weak-mixing with respect to $T\times T$. Hence, the limits of the first three terms in the right-hand side of (2) are 00 in $L^{2}(X\times X,\mu\times\mu)$. Then the theorem follows. ∎
A correspondence principle: Finally, we need the following instance of the correspondence principle, whose proof is classical, and we include it here for sake of completeness.
Theorem 2.14 (Correspondence principle).
Let $G$ be an amenable group, let $A\subset G$ and assume that there exists a left Følner sequence $\Phi$ such that $d_{\Phi}(A)=\lim_{N\to\infty}\frac{|A\cap\Phi_{N}|}{|\Phi_{N}|}$ exists. Then there exist an ergodic $G$-system $(X,\mu,T)$, a clopen set $E\subset X$, a Følner sequence $\Psi$, and a point $a\in\texttt{{gen}}(\mu,\Psi)$ such that $\mu(E)\geq d_{\Phi}(A)$ and $A=\{h\in G:T_{h}a\in E\}$.
Proof.
Consider the compact metric space $X:=\{0,1\}^{G}=\{x=(x_{g})_{g\in G}:x_{g}\in\{0,1\}\>\forall\>g\in G\}$, equipped with the Borel $\sigma$-algebra. We define the continuous action $T$ on $X$ by $T_{h}(x_{g})_{g\in G}=(x_{h^{-1}g})_{g\in G}$, for any $h\in G$ and $(x_{g})_{g\in G}\in X$. Now, consider the point $a=(\mathbbm{1}_{A}(g))_{g\in G}\in X$ and the clopen set $E=\{x=(x_{g})_{g\in G}\in X:x_{e_{G}}=1\}.$ By the choice of $a$, for $h\in G$ we have that $h\in A$ if and only if $T_{h}a\in E$, and therefore $A=\{h\in G:T_{h}a\in E\}.$ Consider the sequence of Borel probability measures on $X$ defined by
$N\mapsto\mu_{N}=\frac{1}{|\Phi_{N}|}\sum_{h\in\Phi_{N}}\delta_{T_{h}a}$
let $\mu^{\prime}$ be a ${}^{\ast}$ limit point of that sequence. Then $\mu^{\prime}(E)=d_{\Phi}(A)$, and $\mu^{\prime}$ is $T$-invariant, but not necessarily ergodic. Let $x\mapsto\mu_{x}^{\prime}$ be the ergodic decomposition of $\mu^{\prime}$. Then $\mu^{\prime}=\int_{x}\mu^{\prime}_{x}~{}\mathrm{d}\mu^{\prime}(x)$, so there exists $x_{0}\in X$ such that for the measure $\mu=\mu_{x_{0}}^{\prime}$, we have that $(X,\mu,T)$ is ergodic and $\mu(E)\geq~{}\mathrm{d}_{\Phi}(A)$. For all $N\in\mathbb{N}$, $\mu_{N}$ is supported on the orbit closure of $a$, and hence $\mu^{\prime}$ is also supported on the orbit closure of $a$. Therefore, we may assume that $\mu$ is also supported on the orbit closure of $a$ (as this is the case with $\mu^{\prime}_{x}$ for $\mu^{\prime}$-almost every $x\in X$). Then it follows by Lemma 2.9, that $a\in\texttt{{gen}}(\mu,\Psi)$ for some Følner sequence $\Psi$. ∎
3. Reduction of Theorem 1.5 to dynamical statements
In this section we translate our first main theorem, namely, Theorem 1.5, in a dynamical language. This will allow us to approach the problem through ergodic theoretic techniques.
3.1. Dynamical reformulation via correspondence principle
Usually in ergodic theory, correspondence principles serve as bridges between combinatorial and dynamical statements. Thus, we can use the correspondence principle (Theorem 2.14) to show that Theorem  1.5 follows from Theorem 3.1 below, which is more dynamical in nature.
Theorem 3.1 (First dynamical reformulation of Theorem 1.5).
Let $G$ be a square absolutely continuous group and $(X,\mu,T)$ be an ergodic $G$-system. Let $a\in\texttt{{gen}}(\mu,\Phi)$ for some Følner sequence $\Phi$ and $E\subset G$ be clopen with $\mu(E)>0$. Then there exist an infinite sequence $B=(b_{n})_{n\in\mathbb{N}}\subset\{h\in G:T_{h}a\in E\}$ and some $t\in G$ such that
[…]
Proof that Theorem 3.1 implies Theorem 1.5.
[…]
3.2. Erdős progressions
The conclusion of Theorem 3.1 is still a rather combinatorial statement, so we need to reformulate it again into a dynamical statement. For this to be achieved, we will use the natural notion of Erdős progressions, as defined in [KMRR23b], which is a dynamical variant of arithmetic progressions.
Definition 3.2.
Given a topological $G$-system $(X,T)$, a point $(x_{0},x_{1},x_{2})\in X^{3}$ is a $3$-term Erdős progression, if there exists an infinite sequence $(g_{n})_{n\in\mathbb{N}}$ in $G$ such that
(3.1) $(T_{g_{n}}\times T_{g_{n}})(x_{0},x_{1})\to(x_{1},x_{2})$
We refer to $3$-term Erdős progressions simply as Erdős progressions. Through the notion of Erdős progressions we are able to reformulate Theorem 3.1 as follows:
Theorem 3.3 (Second dynamical reformulation).
Let $G$ be a square absolutely continuous group, and let $(X,\mu,T)$ be an ergodic $G$-system and $a\in\texttt{{gen}}(\mu,\Phi)$ for some Følner sequence $\Phi$. If $E\subset X$ is a clopen set with $\mu(E)>0$, then there exist $t\in G$, $x_{1}\in E$ and $x_{2}\in{T_{t}}^{-1}E$ such that $(a,x_{1},x_{2})\in X^{3}$ forms an Erdős progression.
For the reduction of Theorem 3.1 to Theorem 3.3 we provide the following lemma.
Lemma 3.4.
[…]
To see how Theorem 3.1 follows from Theorem 3.3 just take $F={T_{t}}^{-1}E$ in the above lemma.
Proof of Lemma 3.4.
By assumption there exists an infinite sequence $(g_{n})_{n\in\mathbb{N}}$ in $G$ such that $(T_{g_{n}}\times T_{g_{n}})(a,x_{1})\to(x_{1},x_{2})$. Since $T_{g_{n}}a\to x_{1}\in E$ and $E$ is open, we get that $T_{g_{n}}a\in E$ for $n$ sufficiently large, so we may assume without loss of generality that $(g_{n})\subset\{g\in G:T_{g}a\in E\}$. Therefore we will construct the sequence $B$ to be a subset of $(g_{n})$.
We construct the sequence $B=(b_{n})_{n\in\mathbb{N}}$ inductively.
  • Since $T_{g_{n}}x_{1}\to x_{2}\in F$ and $F$ is open, we can pick $b_{1}\in(g_{n})$ such that $T_{b_{1}}x_{1}\in F$. Then $(x_{1},x_{2})\in(T^{-1}_{b_{1}}F)\times F$ which is open.
  • Since $(T_{g_{n}}\times T_{g_{n}})(a,x_{1})\to(x_{1},x_{2})$, we pick $b_{2}\in(g_{n})$ such that $(T_{b_{2}}\times T_{b_{2}})(a,x_{1})\in(T^{-1}_{b_{1}}F)\times F$ and $b_{2}\neq b_{1}$ (this is possible since there are infinitely many choices). Then
    $a\in T^{-1}_{{b_{1}}{b_{2}}}F\text{\quad and \quad}x_{1}\in T^{-1}_{b_{2}}F.$
  • Induction step: Assume we have found $b_{1},b_{2},\dots,b_{n}\in(g_{n})$ all distinct to each other such that
    (3.2) $a\in\bigcap_{1\leq i<j\leq n}T^{-1}_{{b_{i}}{b_{j}}}F\text{\quad and \quad}x_{1}\in\bigcap_{1\leq m\leq n}T^{-1}_{b_{m}}F.$
    Since $(T_{g_{n}}\times T_{g_{n}})(a,x_{1})\to(x_{1},x_{2})\in\Big{(}\bigcap_{1\leq m\leq n}T^{-1}_{b_{m}}F\Big{)}\times F$ and this set is open, we can pick $b_{n+1}\in(g_{n})$ such that
    $(T_{b_{n+1}}\times T_{b_{n+1}})(a,x_{1})\in\bigg{(}\bigcap_{1\leq m\leq n}T^{-1}_{b_{m}}F\bigg{)}\times F$
    and $b_{n+1}\not\in\{b_{m}:1\leq m\leq n\}$ (since there are infinitely many choices). Combining this with the inductive hypothesis, we obtain that
    $a\in\bigcap_{1\leq i<j\leq n+1}T^{-1}_{{b_{i}}{b_{j}}}F\text{\quad and \quad}x_{1}\in\bigcap_{1\leq m\leq n+1}T^{-1}_{b_{m}}F.$
[…]
3.3. Continuous factor maps to the Kronecker factor
On our way to show Theorem 3.3, it will be useful to have the extra assumption of the $G$-system having a continuous factor map to its Kronecker factor. The reason for that will become clear towards the proof of our main theorem. As we will see below, it is possible to make such an assumption.
We begin by generalizing a result of Host and Kra in [HK09] from actions of $(\mathbb{N},+)$ to actions of amenable groups.
Lemma 3.5.
[HK09, Proposition 6.1 for group actions] Let $G$ be an amenable group, let $(X,\mu,T)$ be an ergodic $G$-system, $(Z,m,R)$ be its Kronecker factor and $\rho:(X,\mu,T)\to(Z,m,R)$ be a factor map. If $a\in X$ is a transitive point, then there exists a point $z\in Z$ and a Følner sequence $\Psi$ such that
(3.3) $\lim_{N\to\infty}\frac{1}{|\Psi_{N}|}\sum_{g\in\Psi_{N}}f_{1}(T_{g}a)\cdot f_{2}(R_{g}z)=\int_{X}f_{1}\cdot(f_{2}\circ\rho)~{}\mathrm{d}\mu$
holds for any $f_{1}\in C(X)$ and $f_{2}\in C(Z)$.
We remark that the result still holds if we replace $(Z,m,R)$ by any factor of $(X,\mu,T)$ that is distal as a topological system.
Proof.
The proof of this lemma is given in the Appendix A. ∎
Remark 3.6.
Let $(X,\mu,T)$ be an ergodic $G$-system, and let $a\in\texttt{{gen}}(\mu,\Phi)$ for some Følner sequence $\Phi$. From Lemma 2.5 we have that every point in $\texttt{{supp}}(\mu)$ belongs to the orbit closure of $a$, and therefore $\mu\big{(}\overline{\mathcal{O}_{T}(a)}\big{)}=1$. This implies that we can replace $X$ with $\overline{\mathcal{O}_{T}(a)}$ without affecting the ergodic theoretic properties of the system, and then the generic point $a$ is also transitive. Therefore, whenever we have a generic point in a system, we may assume without loss of generality that it is also transitive.
Proposition 3.7.
Let $(X,\mu,T)$ be an ergodic $G$-system, and let $a\in\texttt{{gen}}(\mu,\Phi)$ for some Følner sequence $\Phi$. Then there exists an ergodic extension $(\widetilde{X},\widetilde{\mu},\widetilde{T})$ of $(X,\mu,T)$, a Følner sequence $\widetilde{\Phi}$ and a point $\widetilde{a}\in\texttt{{gen}}(\widetilde{\mu},\widetilde{\Phi})$ such that:
  1. (i)
    There exists a continuous factor map $\widetilde{\pi}:\widetilde{X}\to X$ with $\widetilde{\pi}(\widetilde{a})=a$.
  2. (ii)
    $(\widetilde{X},\widetilde{\mu},\widetilde{T})$ has continuous factor map to its Kronecker factor.
  3. (iii)
    If $(\widetilde{a},\widetilde{x}_{1},\widetilde{x}_{2})\in{\widetilde{X}^{3}}$ is an Erdős progression, then $(a,x_{1},x_{2})\in X^{3}$ is an Erdős progression, where $x_{i}=\widetilde{\pi}(\widetilde{x}_{i})$, for $i=1,2$.
Proof.
The proofs of (i) and (ii) are identical to those in the case $G$ is the semigroup $(\mathbb{N},+)$, and they can be found in [KMRR22, Proposition 3.20]. Therefore, here we only provide a sketch of the proof.
Let $(Z,m,R)$ be the Kronecker factor of $(X,\mu,T)$, and let $\pi:X\to Z$ be a factor map. Define $\widetilde{X}=X\times Z$ and $\widetilde{T}=T\times R$, consider the map $\rho:X\to\widetilde{X}$, given by $\rho(x)=(x,\pi(x))$, and then define $\widetilde{\mu}=\rho\mu$. Then the map $\rho:X\to\widetilde{X}$ is an isomorphism of the $G$-systems $(\widetilde{X},\widetilde{\mu},\widetilde{T})$ and $(X,\mu,T)$, and therefore, since $(X,\mu,T)$ is ergodic, we get that $(\widetilde{X},\widetilde{\mu},\widetilde{T})$ is also ergodic. In addition, the projection on the first coordinate $\widetilde{\pi}:\widetilde{X}\to X$ is a continuous factor map of the systems.
By Remark 3.6, we may assume that the point $a$ is transitive. Then we can use Lemma 3.5 to find a point $z\in Z$ and Følner sequence $\widetilde{\Phi}$ such that (3.3) holds for all $f_{1}\in C(X)$ and $f_{2}\in C(Z)$. Using the definition of the measure $\widetilde{\mu}$, it is not too difficult to see that for any continuous function $F\in C(\widetilde{X})$,
$\lim_{N\to\infty}\frac{1}{|\widetilde{\Phi}_{N}|}\sum_{g\in\widetilde{\Phi}_{N}}F(T_{g}\times R_{g})(a,z)=\int_{\widetilde{X}}F~{}\mathrm{d}\widetilde{\mu},$
which means that the point $\widetilde{a}=(a,z)$ is in $\texttt{{gen}}(\widetilde{\mu},\widetilde{\Phi})$. In addition, $\widetilde{\pi}(\widetilde{a})=a$.
Now, as the systems $(X,\mu,T)$ and $(\widetilde{X},\widetilde{\mu},\widetilde{T})$ are isomorphic, their Kronecker factors are also isomorphic, so we may assume that $(Z,m,R)$ is the Kronecker factor of $(\widetilde{X},\widetilde{\mu},\widetilde{T})$. Then the projection on the second coordinate $p:\widetilde{X}\to Z$ is a continuous factor map from $(\widetilde{X},\widetilde{\mu},\widetilde{T})$ to its Kronecker factor.
Now, let us prove (iii). By assumption, $(\widetilde{T}_{g_{n}}\times\widetilde{T}_{g_{n}})(\widetilde{a},\widetilde{x}_{1})\to(\widetilde{x}_{1},\widetilde{x}_{2})$ for some $(g_{n})_{n\in\mathbb{N}}$ in $G$. We now notice that
$T_{g_{n}}(a)=T_{g_{n}}(\widetilde{\pi}(\widetilde{a}))=\widetilde{\pi}(\widetilde{T}_{g_{n}})\to\widetilde{\pi}(\widetilde{x}_{1})=x_{1},$
since $\widetilde{\pi}$ is a continuous factor map. Similarly, we get $T_{g_{n}}(x_{1})\to x_{2}$ and the result follows. ∎
This proposition allows us to reduce Theorem 3.3 to the case of ergodic $G$-systems with continuous factor maps to their Kronecker factor as desired. Evidently, Theorem 3.3 follows from the following:
Theorem 3.8 (Reduction to $G$-systems with continuous factor maps to the Kronecker factor).
Let $G$ be a square absolutely continuous group, let $(X,\mu,T)$ be an ergodic $G$-system admitting a continuous factor map to its Kronecker factor, and let $a\in\texttt{{gen}}(\mu,\Phi)$ for some Følner sequence $\Phi$. If $E\subset X$ is clopen and $\mu(E)>0$, then there exist $t\in G$, $x_{1}\in E$ and $x_{2}\in T^{-1}_{t}E$ such that $(a,x_{1},x_{2})\in X^{3}$ forms an Erdős progression.
The proof of Theorem 3.8 will be given in the next section.
4. Measures on Erdős progressions and the proof of Theorem 3.8
In this section we prove Theorem 3.8, and consequently, Theorem 1.5.
4.1. Measures on Erdős progressions
In what follows we fix a square absolutely continuous group $G$. We also fix an ergodic $G$-system $(X,\mu,T)$. In addition, as per the assumptions of Theorem 3.8, we assume that $(X,\mu,T)$ admits a continuous factor map to its Kronecker factor. The Kronecker factor of $(X,\mu,T)$ is denoted by $(Z,m,R)$ and $\pi:X\to Z$ stands for the continuous factor map. According to Proposition 2.11, $Z=K/H$ where $K$ is a compact group and $H$ is a closed subgroup of $K$. We denote by $p$ the natural projection $p:K\to K/H$, $p(k)=kH$. We also fix a bi-invariant metric $d_{K}$ on $K$, that is a metric on $K$ compatible with the topology on $K$ such that for all $u,v,w\in K$, $d_{K}(uv,uw)=d_{K}(v,w)=d_{K}(vu,wu)$.
Moreover, $m$ is the (left) Haar measure on $Z$, which is given as the push forward of the (left) Haar measure $m_{K}$ in $K$ by the natural projection $K\to Z=K/H$. We remark that since $K$ is compact, it is unimodular, so $m_{K}$, and consequently $m$, are two-sided invariant. Finally, the action $R=(R_{g})_{g\in G}$ is given by $R_{g}(z)=\alpha(g)z$, where $\alpha:G\to K$ is a continuous group homomorphism with dense image. Also, $\pi_{i}:X\times X\to X$ denotes the projection to the $i$-th coordinate, for $i=1,2$. Moreover, $z\mapsto\eta_{z}$ is a fixed disintegration of $\mu$ over the continuous factor map $\pi$.
Definition 4.1.
Consider the squaring map $s_{K}:K\to K,s_{K}(k)=k^{2}$, on $K$. We define the Borel probability measure $m_{K^{2}}$ on $K$ as the push-forward of the Haar measure $m_{K}$ under the map $s_{K}$, i.e., the measure on $K$, given by $m_{K^{2}}(A)=m_{K}(s_{K}^{-1}(A))$, for each Borel $A\subset K$.
We will prove the following lemma, which is a key ingredient that will allow us to define the measures in order to study Erdős progressions.
Lemma 4.2.
The measure $m_{K^{2}}$ is absolutely continuous with respect to $m_{K}$.
Before we prove Lemma 4.2, let us state and prove an auxiliary lemma that will be used throughout this section:
Lemma 4.3.
Let $\Psi=(\Psi_{N})_{N\in\mathbb{N}}$ be any Følner sequence in $G$. Then the sequence of measures $(\nu_{N})_{N\in\mathbb{N}}$ defined as
$\nu_{N}:=\frac{1}{|\Psi_{N}|}\sum_{g\in\Psi_{N}}\delta_{\alpha(g)}$
converges in the ${}^{\ast}$ topology to the Haar measure $m_{K}$ on $K$.
Proof.
The space of Borel probability measures on $K$ is ${}^{\ast}$ compact and metrizable, so in order to prove the result, it suffices to prove that if $(\nu_{N_{j}})_{j\in\mathbb{N}}$ is a convergent subsequence of $(\nu_{N})_{N\in\mathbb{N}}$, then it converges to the Haar measure $m_{K}$.
Let $(\nu_{N_{j}})_{j\in\mathbb{N}}$ be a subsequence of $(\nu_{N})_{N\in\mathbb{N}}$ which converges in the ${}^{\ast}$ topology to a measure $\nu$ on $K$. To prove that $\nu$ is the Haar measure on $K$, it suffices to prove that for all continuous functions $h$ on $K$ and all $k\in K$, we have that $\int_{K}h(ky)~{}\mathrm{d}\nu(y)=\int_{K}h(y)~{}\mathrm{d}\nu(y)$.
Let $d_{K}$ be a translation invariant metric on $K$. Let also $h:K\to\mathbb{C}$ be continuous, $k\in K$ and $\varepsilon>0$. Since $K$ is compact, we have that $h$ is uniformly continuous, so there is $\delta>0$ such that if $d_{K}(y_{1},y_{2})<\delta$, then $|h(y_{1})-h(y_{2})|<\varepsilon$. Recall that $(\alpha(g))_{g\in G}$ is dense in $K$, so there is $g_{0}$ such that $d_{K}(k,\alpha(g_{0}))<\delta$. We then have that for all $y\in K$, $d_{K}(ky,\alpha(g_{0})y)<\delta$, so $|h(ky)-h(\alpha(g_{0})y)|<\varepsilon$, and we obtain that
$\bigg{|}\int_{K}h(y)~{}\mathrm{d}\nu(y)-\int_{K}h(ky)~{}\mathrm{d}\nu(y)\bigg{|}$ $\leq\bigg{|}\int_{K}h(y)~{}\mathrm{d}\nu(y)-\int_{K}h(\alpha(g_{0})y)~{}\mathrm{d}\nu(y)\bigg{|}+$
$\>\>\>\>\>\>\bigg{|}\int_{K}h(\alpha(g_{0})y)~{}\mathrm{d}\nu(y)-\int_{K}h(ky)~{}\mathrm{d}\nu(y)\bigg{|}$
$\leq\bigg{|}\int_{K}h(y)~{}\mathrm{d}\nu(y)-\int_{K}h(\alpha(g_{0})y)~{}\mathrm{d}\nu(y)\bigg{|}+\varepsilon.$
From the continuity of $h$ and the definition of $\nu$ we have that
$\int_{K}h(\alpha(g_{0})y)~{}\mathrm{d}\nu(y)$ $=\lim_{j\to\infty}\frac{1}{|\Psi_{N_{j}}|}\sum_{g\in\Psi_{N_{j}}}h(\alpha(g_{0})\alpha(g))=\lim_{j\to\infty}\frac{1}{|\Psi_{N_{j}}|}\sum_{g\in\Psi_{N_{j}}}h(\alpha(g_{0}g))$
$=\lim_{j\to\infty}\frac{1}{|\Psi_{N_{j}}|}\sum_{g\in g_{0}\Psi_{N_{j}}}h(\alpha(g))=\lim_{j\to\infty}\frac{1}{|\Psi_{N_{j}}|}\sum_{g\in\Psi_{N_{j}}}h(\alpha(g))$
$=\int_{K}h(y)~{}\mathrm{d}\nu(y),$
so combining with the previous we get that $\big{|}\int_{K}h(y)~{}\mathrm{d}\nu(y)-\int_{K}h(ky)~{}\mathrm{d}\nu(y)\big{|}\leq\varepsilon$, and since $\varepsilon$ was arbitrary, we obtain that $\int_{K}h(y)~{}\mathrm{d}\nu(y)=\int_{K}h(ky)~{}\mathrm{d}\nu(y)$, which proves that $\nu=m_{K}$, and concludes the proof. ∎
Proof of Lemma 4.2.
As $K$ is compact and metrizable, the measures $m_{K},m_{K^{2}}$ are regular. In particular, for each Borel $A$ we have
$m_{K}(A)=\sup_{\begin{subarray}{c}C\subset A\\ C\>\text{compact}\end{subarray}}m_{K}(C)=\inf_{\begin{subarray}{c}O\subset A\\ O\>\text{open}\end{subarray}}m_{K}(O)\>\>\text{ and }\>\>m_{K^{2}}(A)=\sup_{\begin{subarray}{c}C\subset A\\ C\>\text{compact}\end{subarray}}m_{K^{2}}(C)=\inf_{\begin{subarray}{c}O\subset A\\ O\>\text{open}\end{subarray}}m_{K^{2}}(O).$
Therefore, to prove that $m_{K^{2}}$ is absolutely continuous with respect to $m_{K}$, it suffices to prove that for each compact set $C\subset K$, if $m_{K}(C)=0$, then $m_{K^{2}}(C)=0$.
Let $C\subset K$ be a non-empty (for otherwise the result is trivial) compact with $m_{K}(C)=0$ and let $\varepsilon>0$. As $G$ is square absolutely continuous, we know that there are two Følner sequences $\Phi$ and $\Psi$ in $G$ and a $\delta>0$ such that for any $u:G\to[0,1]$ satisfying $\limsup_{N\to\infty}\frac{1}{|\Phi_{N}|}\sum_{g\in\Phi_{N}}u(g)<\delta,$ we have that $\limsup_{N\to\infty}\frac{1}{|\Psi_{N}|}\sum_{g\in\Psi_{N}}u(g^{2})<\varepsilon.$ Since
$0=m_{K}(C)=\inf_{\begin{subarray}{c}O\subset A\\ O\>\text{open}\end{subarray}}m_{K}(O),$
we can pick an open set $O\supset C$ with $m_{K}(O)<\delta$. By Urysohn’s lemma, we know that there is a continuous function $f:K\to[0,1]$ such that $f=1$ on $C$ and $f=0$ outside $O$. By Lemma 4.3 we then have that
$\delta>m_{K}(O)\geq\int_{K}f(k)~{}\mathrm{d}m_{K}(k)=\lim_{N\to\infty}\frac{1}{|\Phi_{N}|}\sum_{g\in\Phi_{N}}f(\alpha(g)),$
and by the choice of $\delta$ we get that
$\limsup_{N\to\infty}\frac{1}{|\Psi_{N}|}\sum_{g\in\Psi_{N}}f(\alpha(g^{2}))<\varepsilon.$
From the definition of $m_{K^{2}}$ and the continuity of $k\mapsto f(k^{2})$ we then obtain
$\int_{K}f(k)~{}\mathrm{d}m_{K^{2}}(k)=\int_{K}f(k^{2})~{}\mathrm{d}m_{K}(k)=\lim_{N\to\infty}\frac{1}{|\Psi_{N}|}\sum_{g\in\Psi_{N}}f(\alpha(g)^{2})=\lim_{N\to\infty}\frac{1}{|\Psi_{N}|}\sum_{g\in\Psi_{N}}f(\alpha(g^{2}))<\varepsilon,$
where for the third equality we use that $\alpha$ is a group homomorphism. So after all we have $m_{K^{2}}(C)\leq\int_{K}f(k)~{}\mathrm{d}m_{K^{2}}(k)<\varepsilon$, and as $\varepsilon$ was arbitrary, we have that $m_{K^{2}}(C)=0$. This concludes the proof. ∎
Now, we define measures $\sigma$ on $X\times X$, and we want $\sigma$ to be defined as natural measures to study Erdős progressions. It is not hard to see that Erdős progressions on $K/H$ are exactly the triplets of the form $(z,kz,k^{2}z)$ for some $k\in K$ and $z\in K/H$. Therefore, following the definition given in [KMRR23b], we will define these measures as the natural measures on points $(x_{1},x_{2})\in X\times X$ to find Erdős progressions on $K/H$ starting at $\pi(a)$, namely $(\pi(a),k\pi(a),k^{2}\pi(a))$, for $k\in K$.
Definition 4.4.
We define the measure $\sigma$ on $X\times X$, given by
(4.1) $\sigma:=\int_{K}\eta_{k\pi(a)}\times\eta_{k^{2}\pi(a)}~{}\mathrm{d}m_{K}(k).$
Let us comment on why the measure $\sigma$ is indeed well-defined. Let $k_{0}\in K$ such that $\pi(a)=k_{0}H$. Since all the maps involved are measurable, we get that the map $K\to M(X)$, $k\mapsto\eta_{k\pi(a)}\times\eta_{k^{2}\pi(a)}~{}\mathrm{d}m_{K}(k)$ is indeed a Borel measurable map on $K$. It remains to prove that it is $m_{K}$ almost everywhere defined, and then the integral will induce a well defined Borel probability measure on $X\times X$.
Since $\eta$ is $m$-almost everywhere well-defined on $Z$, we can consider a set $Z^{\prime}\subset Z$ with $m(Z^{\prime})=1$ such that $\eta_{z}$ is well-defined for all $z\in Z^{\prime}$. Since $pm_{K}=m$, the set $K^{\prime}=p^{-1}(Z^{\prime})$ has $m_{K}(K^{\prime})=1$. Then also $m_{K}(K^{\prime}k_{0}^{-1})=1$, and then it is not difficult to check that the map $K\to M(X)$, $k\mapsto\eta_{k\pi(a)}$ is well defined on $K^{\prime}k_{0}^{-1}$. On the other hand, from Lemma 4.2, we have that $m_{K^{2}}$ is absolutely continuous with respect to $m_{K}$, so $1=m_{K^{2}}(K^{\prime}k_{0}^{-1})=m_{K}(s_{K}^{-1}(K^{\prime}k_{0}^{-1}))$, and again it is not too difficult to check that the map $K\to M(X)$, $k\mapsto\eta_{k^{2}\pi(a)}$ is well defined on $s_{K}^{-1}(K^{\prime}k_{0}^{-1})$.
So, after all, the map $K\to M(X\times X)$, $k\mapsto\eta_{k\pi(a)}\times\eta_{k^{2}\pi(a)}$ is well defined on $K^{\prime}k_{0}^{-1}\cap s_{K}^{-1}(K^{\prime}k_{0}^{-1})$ and $m_{K}(K^{\prime}k_{0}^{-1}\cap s_{K}^{-1}(K^{\prime}k_{0}^{-1}))=1$, i.e., $k\mapsto\eta_{k\pi(a)}\times\eta_{k^{2}\pi(a)}$ is $m_{K}$-almost everywhere well-defined, which was to be proved, and therefore $\sigma$ is indeed well-defined.
Using the invariance of $m_{K}$ we can express $\sigma$ as
(4.2) $\sigma=\int_{K}\eta_{kk_{0}H}\times\eta_{k^{2}k_{0}H}~{}\mathrm{d}m_{K}(k)=\int_{K}\eta_{kH}\times\eta_{kk_{0}^{-1}kH}~{}\mathrm{d}m_{K}(k)$
Proposition 4.5.
The measure $\sigma$ has the following properties:
  1. (i)
    $\pi_{1}\sigma=\mu$.
  2. (ii)
    $\pi_{2}\sigma$ is absolutely continuous with respect to $\mu$.
Proof of Proposition 4.5.
(i) Using (4.2), we have that
$\pi_{1}\sigma=\int_{K}\eta_{kH}~{}\mathrm{d}m_{K}(k)=\int_{Z}\eta_{z}~{}\mathrm{d}m(z)=\mu.$
(ii) From the definition of $\sigma$ we have that $\pi_{2}\sigma=\int_{K}\eta_{k^{2}\pi(a)}~{}\mathrm{d}m_{K}(k).$ Fix $k_{0}$ such that $\pi(a)=k_{0}H$.
Let $A\subset X$ with $\mu(A)=0$. Then $\mu(A)=\int_{Z}\eta_{z}(A)~{}\mathrm{d}m(z)$, so we get that there is a set $Z^{\prime}\subset Z$ with $m(Z^{\prime})=1$ such that for all $z\in Z^{\prime}$, $\eta_{z}(A)=0$. As $pm_{K}=m$, we have $m_{K}(p^{-1}(Z^{\prime}))=1$ and then also $m_{K}((p^{-1}Z^{\prime})k_{0}^{-1})=1$. Finally, using Lemma 4.2 we get that $m_{K^{2}}((p^{-1}Z^{\prime})k_{0}^{-1})=m_{K}(s_{K}^{-1}((p^{-1}Z^{\prime})k_{0}^{-1}))=1$. For each $k\in s_{K}^{-1}((p^{-1}Z^{\prime})k_{0}^{-1})$ we have that $k^{2}\pi(a)\in Z^{\prime}$, so $\eta_{k^{2}\pi(a)}(A)=0$, and therefore $\pi_{2}\sigma(A)=0$. ∎
Theorem 4.6.
For any set $E\subset X$ with $\mu(E)>0$, we have that
$\sigma\bigg{(}E\times\bigcup_{t\in G}T^{-1}_{t}E\bigg{)}>0.$
Proof.
Let $E\subset X$ with $\mu(E)>0$ and recall that we want to show that the set $E\times\bigcup_{t\in G}T^{-1}_{t}E$ has positive measure $\sigma$. We begin by expressing this set as
$E\times\bigcup_{t\in G}T^{-1}_{t}E=(E\times X)\cap\bigg{(}X\times\bigcup_{t\in G}T^{-1}_{t}E\bigg{)}.$
By Proposition 4.5 (i), we have that $\sigma(E\times X)=\mu(E)>0$. Therefore, it is enough to show that
(4.3) $\sigma\bigg{(}X\times\bigcup_{t\in G}T^{-1}_{t}E\bigg{)}=1.$
Notice that the set $\bigcup_{t\in G}T_{t}^{-1}E$ is clearly $T$-invariant and since $\mu$ is ergodic and $\mu(E)>0$, it follows that $\mu\big{(}\bigcup_{t\in G}T_{t}^{-1}E\big{)}=1$. By Proposition 4.5 (ii), $\pi_{2}\sigma$ is absolutely continuous with respect to $\mu$, so
$1=\pi_{2}\sigma\bigg{(}\bigcup_{t\in G}T_{t}^{-1}E\bigg{)}=\sigma\bigg{(}X\times\bigcup_{t\in G}T_{t}^{-1}E\bigg{)},$
which concludes the proof. ∎
4.2. A continuous ergodic decomposition
In this subsection we will define measures $\lambda_{(x_{1},x_{2})}$, for $(x_{1},x_{2})\in X\times X$, in a way that $(x_{1},x_{2})\mapsto\lambda_{(x_{1},x_{2})}$ will be a continuous ergodic decomposition of $\mu\times\mu$. We follow the definition given in [KMRR22, Eq. (3.10)] and [KMRR23b, Eq. (3.1)].
Definition 4.7.
For $(x_{1},x_{2})\in X\times X$ we define the measures $\lambda_{(x_{1},x_{2})}$ on $X\times X$ by
(4.4) $\lambda_{(x_{1},x_{2})}=\int_{K}\eta_{k\pi(x_{1})}\times\eta_{k\pi(x_{2})}~{}\mathrm{d}m_{K}(k).$
Given $x_{1},x_{2}\in X$, we let $k_{1},k_{2}\in K$ be such that $\pi(x_{i})=k_{i}H$, for $i=1,2$. Then, using the invariance of $m_{K}$, we can write
(4.5) $\lambda_{(x_{1},x_{2})}=\int_{K}\eta_{kk_{1}H}\times\eta_{kk_{2}H}~{}\mathrm{d}m_{K}(k)=\int_{K}\eta_{kH}\times\eta_{kk_{1}^{-1}k_{2}H}~{}\mathrm{d}m_{K}(k).$
Theorem 4.8.
The map $(x_{1},x_{2})\mapsto\lambda_{(x_{1},x_{2})}$ is a continuous ergodic decomposition of $\mu\times\mu$ in the following sense:
  • (i)
    It is a continuous map.
  • (ii)
    It satisfies $\int_{X\times X}\lambda_{(x_{1},x_{2})}~{}\mathrm{d}(\mu\times\mu)(x_{1},x_{2})=\mu\times\mu$.
  • (iii)
    The $G$-system $(X\times X,\lambda_{(x_{1},x_{2})},T\times T)$ is ergodic for $\mu\times\mu$-almost every $(x_{1},x_{2})\in X\times X$.
In addition, for any $x_{1},x_{2}\in X$ we have that
(4.6) $\lambda_{(x_{1},x_{2})}=\lambda_{(T_{g}x_{1},T_{g}x_{2})}$
for any $g\in G$.
Proof.
For the proof of (i) and (ii) we refer to [KMRR22, Proposition 3.11], as the proof there can be directly adapted to our case. We will now prove (iii). It is not too difficult to see that for all $(x_{1},x_{2})\in X\times X$ and for all $g\in G$, $(T_{g}\times T_{g})\lambda_{(x_{1},x_{2})}=\lambda_{(x_{1},x_{2})}$, i.e., $\lambda_{(x_{1},x_{2})}$ is $T\times T$-invariant. Therefore, to prove (iii), it suffices to prove that there is some Følner sequence $\Psi$ in $G$ such that for $(\mu\times\mu)$-almost every $(x_{1},x_{2})\in X\times X$ and all bounded and measurable functions $F$ on $X\times X$,
$\lim_{N\to\infty}\frac{1}{|\Psi_{N}|}\sum_{g\in\Psi_{N}}(T_{g}\times T_{g})F=\int_{X\times X}F~{}\mathrm{d}\lambda_{(x_{1},x_{2})},$
in $L^{2}(X\times X,\lambda_{(x_{1},x_{2})})$.
Now, since $X$ is a compact metric space, there is a countable family of continuous functions $(f_{k})_{k\in\mathbb{N}}$ which is dense in $L^{p}(\nu)$ for all $p\in[1,+\infty)$ and all Borel probability measures $\nu$ on $X$. Then, it is not too difficult to see that the set consisting of finite linear combinations of functions the form $(f_{j_{1}}\otimes f_{j_{2}})_{j_{1},j_{2}\in\mathbb{N}}$ is dense in $L^{2}(\rho)$ for all Borel probability measures $\rho$ on $X\times X$. Hence, using an approximation argument, it suffices to prove that there is a Følner $\Psi$ in $G$ and a set $W\subset X\times X$ with $(\mu\times\mu)(W)=1$ such that for all $(x_{1},x_{2})\in W$ and for all $j_{1},j_{2}\in\mathbb{N}$
$\lim_{N\to\infty}\frac{1}{|\Psi_{N}|}\sum_{g\in\Psi_{N}}(T_{g}\times T_{g})(f_{j_{1}}\otimes f_{j_{2}})=\int_{X\times X}f_{j_{1}}\otimes f_{j_{2}}~{}\mathrm{d}\lambda_{(x_{1},x_{2})},$
in $L^{2}(X\times X,\lambda_{(x_{1},x_{2})})$.
Step 1. Let $\Phi$ be any Følner sequence in $G$. Then, using Theorem 2.12, we get that for each $j_{1},j_{2}\in\mathbb{N}$,
$\lim_{N\to\infty}\frac{1}{|\Phi_{N}|}\sum_{g\in\Phi_{N}}(T_{g}\times T_{g})(f_{j_{1}}\otimes f_{j_{2}})=\lim_{N\to\infty}\frac{1}{|\Phi_{N}|}\sum_{g\in\Phi_{N}}(T_{g}\times T_{g})(\mathbb{E}_{\mu}(f_{j_{1}}\>|\>Z)\otimes\mathbb{E}_{\mu}(f_{j_{2}}\>|\>Z))$
in $L^{2}(\mu\times\mu)$. Combining this with (ii) yields
$\lim_{N\to\infty}\int_{X\times X}\int_{X\times X}\bigg{|}\frac{1}{|\Phi_{N}|}\sum_{g\in\Phi_{N}}(T_{g}\times T_{g})(f_{j_{1}}\otimes f_{j_{2}})(y_{1},y_{2})$
$\>\>\>\>\>\>-\frac{1}{|\Phi_{N}|}\sum_{g\in\Phi_{N}}(T_{g}\times T_{g})(\mathbb{E}_{\mu}(f_{j_{1}}\>|\>Z)\otimes\mathbb{E}_{\mu}(f_{j_{2}}\>|\>Z))(y_{1},y_{2})\bigg{|}^{2}~{}\mathrm{d}\lambda_{(x_{1},x_{2})}(y_{1},y_{2})~{}\mathrm{d}(\mu\times\mu)(x_{1},x_{2})=0.$
Then for each $j_{1},j_{2}\in\mathbb{N}$ we can find a sub-Følner sequence $\widetilde{\Phi}$ of $\Phi$, depending on $j_{1},j_{2}$, such that for $(\mu\times\mu)$-almost every $(x_{1},x_{2})\in X\times X$, we have
$\lim_{N\to\infty}\int_{X\times X}\bigg{|}\frac{1}{|\widetilde{\Phi}_{N}|}\sum_{g\in\widetilde{\Phi}_{N}}(T_{g}\times T_{g})(f_{j_{1}}\otimes f_{j_{2}})$
$\hskip 162.18062pt-\frac{1}{|\widetilde{\Phi}_{N}|}\sum_{g\in\widetilde{\Phi}_{N}}(T_{g}\times T_{g})(\mathbb{E}_{\mu}(f_{j_{1}}\>|\>Z)\otimes\mathbb{E}_{\mu}(f_{j_{2}}\>|\>Z))\bigg{|}^{2}~{}\mathrm{d}\lambda_{(x_{1},x_{2})}=0,$
and since the limits of both averages above exist by Theorem 2.1, we have that, for $(\mu\times\mu)$-almost every $(x_{1},x_{2})\in X\times X$,
$\lim_{N\to\infty}\frac{1}{|\widetilde{\Phi}_{N}|}\sum_{g\in\widetilde{\Phi}_{N}}(T_{g}\times T_{g})(f_{j_{1}}\otimes f_{j_{2}})=\lim_{N\to\infty}\frac{1}{|\widetilde{\Phi}_{N}|}\sum_{g\in\widetilde{\Phi}_{N}}(T_{g}\times T_{g})(\mathbb{E}_{\mu}(f_{j_{1}}\>|\>Z)\otimes\mathbb{E}_{\mu}(f_{j_{2}}\>|\>Z))$
in $L^{2}(X\times X,\lambda_{(x_{1},x_{2})})$. Since the family $(f_{j_{1}}\otimes f_{j_{2}})_{j_{1},j_{2}\in\mathbb{N}}$ is countable, then using a diagonal argument one can find a Følner sequence $\Psi$ and a set $W_{1}\subset X\times X$ with $(\mu\times\mu)(W_{1})=1$ such that for all $(x_{1},x_{2})\in W_{1}$ and $j_{1},j_{2}\in\mathbb{N}$, we have that
(4.7) $\lim_{N\to\infty}\frac{1}{|\Psi_{N}|}\sum_{g\in\Psi_{N}}(T_{g}\times T_{g})(f_{j_{1}}\otimes f_{j_{2}})=\lim_{N\to\infty}\frac{1}{|\Psi_{N}|}\sum_{g\in\Psi_{N}}(T_{g}\times T_{g})(\mathbb{E}_{\mu}(f_{j_{1}}\>|\>Z)\otimes\mathbb{E}_{\mu}(f_{j_{2}}\>|\>Z))$
in $L^{2}(X\times X,\lambda_{(x_{1},x_{2})})$.
Step 2. Consider the sequence of probability measures on $K$ defined by $\nu_{N}:=\frac{1}{|\Psi_{N}|}\sum_{g\in\Psi_{N}}\delta_{\alpha(g)}$. From Lemma 4.3 we know that $\nu_{N}\to m_{K}$ as $N\to\infty$ in the ${}^{\ast}$ topology.
Let $\phi_{1},\phi_{2}:Z\to\mathbb{C}$ be continuous. For each $z_{1},z_{2}\in Z$, consider the function $\phi_{z_{1},z_{2}}:K\to\mathbb{C}$ defined by $\phi_{z_{1},z_{2}}(k)=\phi_{1}(kz_{1})\phi_{2}(kz_{2})$. Then, $\phi_{z_{1},z_{2}}$ is continuous, so we have that
$\int_{K}\phi_{1}(kz_{1})\phi_{2}(kz_{2})~{}\mathrm{d}m_{K}(k)$ $=\int_{K}\phi_{z_{1},z_{2}}(k)~{}\mathrm{d}m_{K}(k)=\lim_{N\to\infty}\frac{1}{|\Psi_{N}|}\sum_{g\in\Psi_{N}}\phi_{z_{1},z_{2}}(\alpha(g))$
$=\lim_{N\to\infty}\frac{1}{|\Psi_{N}|}\sum_{g\in\Psi_{N}}\phi_{1}(\alpha(g)z_{1})\phi_{2}(\alpha(g)z_{2})$
$=\lim_{N\to\infty}\frac{1}{|\Psi_{N}|}\sum_{g\in\Psi_{N}}(R_{g}\times R_{g})(\phi_{1}\otimes\phi_{2})(z_{1},z_{2}).$
Since the previous holds for all $z_{1},z_{2}\in Z$, using the dominated convergence theorem, we get that
$\lim_{N\to\infty}\int_{Z\times Z}\bigg{|}\frac{1}{|\Psi_{N}|}\sum_{g\in\Psi_{N}}(R_{g}\times R_{g})(\phi_{1}\otimes\phi_{2})(z_{1},z_{2})-\int_{K}\phi_{1}(kz_{1})\phi_{2}(kz_{2})~{}\mathrm{d}m_{K}(k)\bigg{|}^{2}~{}\mathrm{d}(m\times m)(z_{1},z_{2})=0,$
i.e., the sequence $\frac{1}{|\Psi_{N}|}\sum_{g\in\Psi_{N}}(R_{g}\times R_{g})(\phi_{1}\otimes\phi_{2})$ converges in $L^{2}(Z\times Z,m\times m)$ as $N\to\infty$ to the function $(z_{1},z_{2})\mapsto\int_{K}\phi_{1}(kz_{1})\phi_{2}(kz_{2})~{}\mathrm{d}m_{K}(k)$.
Step 3. Given two bounded measurable maps $h_{1},h_{2}:Z\to\mathbb{C}$, approximating them in $L^{2}(Z,m)$ by two continuous functions $\phi_{1},\phi_{2}$ and using Step 2, one can prove that $\frac{1}{|\Psi_{N}|}\sum_{g\in\Psi_{N}}(R_{g}\times R_{g})(h_{1}\otimes h_{2})$ converges in $L^{2}(Z\times Z,m\times m)$ as $N\to\infty$ to the function $(z_{1},z_{2})\mapsto\int_{K}h_{1}(kz_{1})h_{2}(kz_{2})~{}\mathrm{d}m_{K}(k)$. Since $\pi:(X,\mu,T)\to(Z,m,R)$ is a factor map, it is not too difficult then to see that
(4.8) $\frac{1}{|\Psi_{N}|}\sum_{g\in\Psi_{N}}(T_{g}\times T_{g})(h_{1}\circ\pi\otimes h_{2}\circ\pi)\to\Big{[}(x_{1},x_{2})\mapsto\int_{K}h_{1}(k\pi(x_{1}))h_{2}(k\pi(x_{2}))~{}\mathrm{d}m_{K}(k)\Big{]}$
as $N\to\infty$ in $L^{2}(X\times X,\mu\times\mu)$.
For each $j\in\mathbb{N}$, $\mathbb{E}_{\mu}(f_{j}\>|\>Z)$ can be viewed either as a function on $Z$ or as a function on $X$ measurable with respect to $\pi^{-1}(Z)$. In this proof, we always view $\mathbb{E}_{\mu}(f_{j}\>|\>Z)$ as a function on $X$ measurable with respect to $\pi^{-1}(Z)$. For each $j\in\mathbb{N}$, let $\psi_{j}$ be $\mathbb{E}_{\mu}(f_{j}\>|\>Z)$ when viewed as a function on $Z$, so we have that for $\mu$-almost every $x\in X$, $\psi_{j}\circ\pi(x)=\mathbb{E}_{\mu}(f_{j}\>|\>Z)(x)$. Then for each $j_{1},j_{2}\in\mathbb{N}$ and $(x_{1},x_{2})\in X\times X$, we have
$\int_{X\times X}f_{j_{1}}(y_{1})f_{j_{2}}(y_{2})~{}\mathrm{d}\lambda_{(x_{1},x_{2})}(y_{1},y_{2})$ $=\int_{K}\int_{X\times X}f_{j_{1}}(y_{1})f_{j_{2}}(y_{2})~{}\mathrm{d}(\eta_{k\pi(x_{1})}\times\eta_{k\pi(x_{2})})(y_{1},y_{2})~{}\mathrm{d}m_{K}(k)$
$=\int_{K}\int_{X}f_{j_{1}}(y_{1})~{}\mathrm{d}\eta_{k\pi(x_{1})}(y_{1})\int_{X}f_{j_{2}}(y_{2})~{}\mathrm{d}\eta_{k\pi(x_{2})}(y_{2})~{}\mathrm{d}m_{K}(k)$
(4.9) $=\int_{K}\psi_{j_{1}}(k\pi(x_{1}))\psi_{j_{1}}(k\pi(x_{1}))~{}\mathrm{d}m_{K}(k)$
where the last equality follows using (2.1). After all, combining (4.8) and (4.2) we get that for each $j_{1},j_{2}\in\mathbb{N}$,
$\frac{1}{|\Psi_{N}|}\sum_{g\in\Psi_{N}}(T_{g}\times T_{g})(\mathbb{E}_{\mu}(f_{j_{1}}\>|\>Z)\otimes\mathbb{E}_{\mu}(f_{j_{2}}\>|\>Z))\to\Big{[}(x_{1},x_{2})\mapsto\int_{X\times X}f_{j_{1}}\otimes f_{j_{2}}~{}\mathrm{d}\lambda_{(x_{1},x_{2})}\Big{]}$
as $N\to\infty$ in $L^{2}(X\times X,\mu\times\mu)$. Now, since the family $f_{j_{1}}\otimes f_{j_{2}}$ is countable, using (ii) and a diagonal argument as in Step 1, one can find a sub-Følner sequence of $\Psi$, which by abuse of notation we again denote by $\Psi$, and a set $W_{2}\subset X\times X$ with $(\mu\times\mu)(W_{2})=1$ such that for all $(x_{1},x_{2})\in W_{2}$ and $j_{1},j_{2}\in\mathbb{N}$,
(4.10) $\lim_{N\to\infty}\frac{1}{|\Psi_{N}|}\sum_{g\in\Psi_{N}}(T_{g}\times T_{g})(\mathbb{E}_{\mu}(f_{j_{1}}\>|\>Z)\otimes\mathbb{E}_{\mu}(f_{j_{2}}\>|\>Z))=\int_{X\times X}f_{j_{1}}\otimes f_{j_{2}}~{}\mathrm{d}\lambda_{(x_{1},x_{2})}$
in $L^{2}(X\times X,\lambda_{(x_{1},x_{2})})$.
Let $W=W_{1}\cap W_{2}$. Then $(\mu\times\mu)(W)=1$, and combining (4.7) and (4.10), we get that for all $(x_{1},x_{2})\in W$ and all $j_{1},j_{2}\in\mathbb{N}$,
$\lim_{N\to\infty}\frac{1}{|\Psi_{N}|}\sum_{g\in\Psi_{N}}(T_{g}\times T_{g})(f_{j_{1}}\otimes f_{j_{2}})=\int_{X\times X}f_{j_{1}}\otimes f_{j_{2}}~{}\mathrm{d}\lambda_{(x_{1},x_{2})}$
in $L^{2}(X\times X,\lambda_{(x_{1},x_{2})})$, which was to be proved.
To conclude the proof we are left with showing (4.6). To this end, using the invariance of $m_{K}$, for any $g\in G$, we have that
$\lambda_{(T_{g}x_{1},T_{g}x_{2})}$ $=\int_{Z}\eta_{k\alpha(g)\pi(x_{1})}\times\eta_{k\alpha(g)\pi(x_{2})}~{}\mathrm{d}m_{K}(k)=\int_{Z}\eta_{k\pi(x_{1})}\times\eta_{k\pi(x_{2})}~{}\mathrm{d}m_{K}(k\alpha(g)^{-1})$
$=\int_{Z}\eta_{k\pi(x_{1})}\times\eta_{k\pi(x_{2})}~{}\mathrm{d}m_{K}(k)=\lambda_{(x_{1},x_{2})}.$
The proof of the theorem is now complete. ∎
Theorem 4.9.
We have that
$\sigma(\{(x_{1},x_{2})\in X\times X:(x_{1},x_{2})\in\texttt{{supp}}(\lambda_{(x_{1},x_{2})})\})=1.$
Theorem 4.10.
There exists some Følner sequence $\Psi$ such that
$\sigma(\{(x_{1},x_{2})\in X\times X:(a,x_{1})\in\texttt{{gen}}(\lambda_{(x_{1},x_{2})},\Psi)\})=1.$
We first deal with Theorem 4.9. Let us first state and prove some results that will be useful in order to prove Theorem 4.9.
Let $\mathcal{F}(X)$ be the family of the closed, nonempty subsets of the compact metric space $(X,d_{X})$. We endow this family with the Hausdorff metric D, defined by
$\texttt{{D}}(A,B)=\max\bigg{\{}\sup_{x\in A}d_{X}(x,B),\sup_{y\in B}d_{X}(y,A)\bigg{\}},$
for any $A,B\in\mathcal{F}(X)$.
We will need the following two lemmas, the proofs of which are omitted, as they can be found in [KMRR23b]:
Lemma 4.11.
[KMRR23b, Lemma 3.8] Let $W$ be a compact metric space, $M(W)$ the space of Borel probability measures on $W$ endowed with the ${}^{\ast}$ topology, and $\mathcal{F}(W)$ the space of closed, non-empty subsets of $W$ with the Hausdorff metric. Then
  • The map $\nu\mapsto\texttt{{supp}}(\nu)$ from $M(W)$ to $\mathcal{F}(W)$ is Borel measurable.
  • If $x\mapsto\rho_{x}$ is a measurable map from $W$ to $M(W)$, then $\{x\in X:x\in\texttt{{supp}}(\rho_{x})\}$ is a Borel set.
Lemma 4.12.
[KMRR23b, Lemma 3.9] The disintegration $z\mapsto\eta_{z}$ satisfies that the subset
$\{x\in X:x\in\texttt{{supp}}(\eta_{x})\}$ of $X$ is Borel measurable and
$\mu(\{x\in X:x\in\texttt{{supp}}(\eta_{x})\})=1.$
Using those, we can now prove the following proposition, which is a variant of [KMRR23b, Proposition 3.10].
Proposition 4.13.
There exists a sequence $\delta_{j}\to 0$ such that for $\mu$-almost every $x\in X$ there exists $w\in K$ with $\pi(x)=wH$ such that for any open neighborhood $U$ of $x$, we have
(4.11) $\lim_{j\to\infty}\frac{m_{K}(\{k\in K\colon\eta_{kH}(U)>0\}\cap\texttt{{B}}_{K}(w,\delta_{j}))}{m_{K}(\texttt{{B}}_{K}(w,\delta_{j}))}=1,$
where $\texttt{{B}}_{K}(w,\delta_{j})$ denotes the ball centered at $w\in K$ and with radius $\delta_{j}$ in $K$ with respect to the fixed metric $d_{K}$.
Proof.
Let $F:K\to\mathcal{F}(X)$ given by $F(k)=\texttt{{supp}}(\eta_{kH})$. The natural projection $p:K\to Z$ is continuous, hence Borel measurable, the map $z\mapsto\eta_{z}$ is Borel measurable, as $(\eta_{z})_{z\in Z}$ is a disintegration and by Lemma 4.11, the map $\nu\mapsto\texttt{{supp}}(\nu)$ is also Borel measurable, thus, we obtain that their composition $F$ is also Borel measurable. By Lusin’s theorem [AB06, Theorem 12.8], for any $j\in\mathbb{N}$ there exists a closed $K_{j}\subset K$ with
(4.12) $m_{K}(K_{j})>1-2^{-j}$
such that $F|_{K_{j}}$ is continuous. For $j\in\mathbb{N}$, using the fact that $K$, and so $K_{j}$, is compact, we obtain that $F|_{K_{j}}$ is uniformly continuous. Therefore, for any $j\in\mathbb{N}$, there exists some $\delta_{j}>0$ such that for any $k_{1},k_{2}\in K_{j}$ we have
$d_{K}(k_{1},k_{2})\leq\delta_{j}\Longrightarrow\texttt{{D}}(F(k_{1}),F(k_{2}))<\frac{1}{j}.$
Fix $j\in\mathbb{N}$. Then by the invariance of $m_{K}$, there is $c_{j}>0$ such that for all $k\in K$, $m_{K}(\texttt{{B}}_{K}(k,\delta_{j}))=c_{j}$. The regularity of the measure $m_{K}$ implies that there is a compact set $C_{j}\subset\texttt{{B}}_{K}(e_{K},\delta_{j})$ such that $m_{K}(C_{j})\geq c_{j}-\frac{c_{j}}{2^{j}}$. Now, by Urysohn’s lemma, there is a continuous function $f_{j}:K\to[0,1]$ such that $f_{j}=1$ on $C_{j}$ and $f_{j}=0$ outside $\texttt{{B}}_{K}(e_{K},\delta_{j})$.
Consider the set
$W_{j}=\bigg{\{}k\in K_{j}:\int_{K_{j}}f_{j}(wk^{-1})~{}\mathrm{d}m_{K}(w)\geq\bigg{(}1-\frac{1}{j}\bigg{)}\int_{K}f_{j}(w)~{}\mathrm{d}m_{K}(w)\bigg{\}}.$
Note that
$\int_{K}f_{j}(w)~{}\mathrm{d}m_{K}(w)\geq m_{K}(C_{j})\geq c_{j}-\frac{c_{j}}{2^{j}}>0.$
Then we can consider the function
$\chi_{j}:K\to[0,1],\>\>\chi_{j}(k)=\frac{\int_{K_{j}}f_{j}(wk^{-1})~{}\mathrm{d}m_{K}(w)}{\int_{K}f_{j}(w)~{}\mathrm{d}m_{K}(w)}$
and moreover, we let
$A_{j}=\bigg{\{}k\in K:\chi_{j}(k)\geq 1-\frac{1}{j}\bigg{\}}.$
Then we see that
(4.13) $W_{j}=K_{j}\cap A_{j}.$
We will show that the set $W_{j}$ is closed. Using the dominated convergence theorem and the fact that $f_{j}$ is continuous, one can show that if $(k_{\ell})_{\ell\in\mathbb{N}}$ is a sequence in $K$ and $k_{\ell}\to k$, then $\int_{K_{j}}f_{j}(wk_{\ell}^{-1})~{}\mathrm{d}m_{K}(w)\to\int_{K_{j}}f_{j}(wk^{-1})~{}\mathrm{d}m_{K}(w)$, and this proves the continuity of $\chi_{j}$. As a result, the set $A_{j}$ is a closed subset of $K$, and since $K_{j}$ is also closed, it follows that $W_{j}$ is closed.
Now, using Fubini’s theorem and the invariance of $m_{K}$ we deduce that
$\int_{K}\chi_{j}(k)~{}\mathrm{d}m_{K}(k)$ $=\frac{1}{\int_{K}f_{j}(w)~{}\mathrm{d}m_{K}(w)}\int_{K}\int_{K}f_{j}(wk^{-1})\mathbbm{1}_{K_{j}}(w)~{}\mathrm{d}m_{K}(k)~{}\mathrm{d}m_{K}(w)$
$=\frac{1}{\int_{K}f_{j}(w)~{}\mathrm{d}m_{K}(w)}\int_{K}\mathbbm{1}_{K_{j}}(w)\int_{K}f_{j}(k)~{}\mathrm{d}m_{K}(k)~{}\mathrm{d}m_{K}(w)$
$=m_{K}(K_{j})>1-\frac{1}{2^{j}}$
and then we have that
$1-\frac{1}{2^{j}}<\int_{A_{j}}\chi_{j}(k)~{}\mathrm{d}m_{K}(k)+\int_{A_{j}^{\text{c}}}\chi_{j}(k)~{}\mathrm{d}m_{K}(k)\leq m_{K}(A_{j})+\bigg{(}1-\frac{1}{j}\bigg{)}m_{K}(A_{j}^{\text{c}})=1-\frac{1}{j}+\frac{m_{K}(A_{j})}{j},$
which gives that
(4.14) $m_{K}(A_{j})>1-\frac{j}{2^{j}}.$
Combining (4.12), (4.13) and (4.14), we obtain that
$\sum_{j\in\mathbb{N}}m_{K}(K\setminus W_{j})=\sum_{j\in\mathbb{N}}m_{K}(K\setminus(K_{j}\cap A_{j}))<\sum_{j\in\mathbb{N}}\frac{1+j}{2^{j}}<\infty.$
Let $W=\bigcup_{J\in\mathbb{N}}\bigcap_{j\geq J}W_{j}$. It follows by the Borel- Cantelli lemma, using the last equation above, that $m_{K}(W)=1$. Now let $L:=\{x\in X\colon x\in\texttt{{supp}}(\eta_{\pi(x)})\}\cap\pi^{-1}(p(W))$. Observe that $p(W)=\bigcup_{J\in\mathbb{N}}p\big{(}\bigcap_{j\geq J}W_{j}\big{)}$. For each $J\in\mathbb{N}$, $\bigcap_{j\geq J}W_{j}$ is a closed subset of $K$, thus it is compact and since $p$ is continuous, we get that $p\big{(}\bigcap_{j\geq J}W_{j}\big{)}$ is also compact, thus it is Borel measurable. As a result, we get that $p(W)$ is indeed a Borel subset of $Z$. In addition, $p^{-1}(p(W))\supset W$ and since $m_{K}(W)=1$ we have that $m_{K}(p^{-1}(p(W))=1$. Therefore, $m(p(W))=1$ and hence $\mu(\pi^{-1}(p(W)))=1$. Then, in view of Lemma 4.12, it follows that $L$ is a Borel subset of $X$ and $\mu(L)=1$.
We now show that elements of $L$ satisfy (4.11) and this will conclude the proof. Let $x\in L=\{x\in X\colon x\in\texttt{{supp}}(\eta_{\pi(x)})\}\cap\pi^{-1}(p(W))$. Then $x\in\pi^{-1}(p(W))$ so there is $w\in W$ such that $\pi(x)=p(w)=wH$. Let $U$ be an open neighborhood of $x$. Then we have that there exists $J\in\mathbb{N}$ such that for any $j\geq J$, $w\in W_{j}$ and $\texttt{{B}}(x,\tfrac{1}{j})\subset U$.
We now claim that
$\texttt{{B}}_{K}(w,\delta_{j})\cap K_{j}\subset\{k\in K\colon\eta_{kH}(U)>0\}.$
To prove this, we let $w^{\prime}\in\texttt{{B}}_{K}(w,\delta_{j})\cap K_{j}$. Then $d_{K}(w^{\prime},w)<\delta_{j}$ and so, $\texttt{{D}}(F(w^{\prime}),F(w))<\tfrac{1}{j}$. Notice now that $F(w)=\texttt{{supp}}(\eta_{wH})=\texttt{{supp}}(\eta_{\pi(x)})$, and so, $x\in F(w)$, as $x\in L$. Then, by the definition of the Hausdorff metric, there exists $x^{\prime}\in F(w^{\prime})$ with $d_{X}(x,x^{\prime})<\tfrac{1}{j}$, and so, $x^{\prime}\in U$, which, combined with the fact that $x^{\prime}\in F(w^{\prime})$, yields $U\cap F(w^{\prime})\neq\emptyset$. It follows that $\eta_{w^{\prime}H}(U)>0$.
It follows from the above claim that
(4.15) $\frac{m_{K}(\{k\in K\colon\eta_{kH}(U)>0\}\cap\texttt{{B}}_{K}(w,\delta_{j}))}{m_{K}(\texttt{{B}}_{K}(w,\delta_{j}))}\geq\frac{m_{K}(\texttt{{B}}_{K}(w,\delta_{j})\cap K_{j})}{m_{K}(\texttt{{B}}_{K}(w,\delta_{j}))}=\frac{\int_{K}\mathbbm{1}_{\texttt{{B}}_{K}(w,\delta_{j})}(u)\mathbbm{1}_{K_{j}}(u)~{}\mathrm{d}m_{K}(u)}{m_{K}(\texttt{{B}}_{K}(e_{K},\delta_{j}))}$
The denominator in the right-most term in (4.15) is smaller or equal to $\frac{2^{j}}{2^{j}-1}m_{K}(C_{j})$, which then is smaller or equal to $\frac{2^{j}}{2^{j}-1}\int_{K}f_{j}(u)~{}\mathrm{d}m_{K}(u)$, and therefore the expression in (4.15) is greater or equal to
(4.16) $\frac{\int_{K}\mathbbm{1}_{\texttt{{B}}_{K}(w,\delta_{j})}(u)\mathbbm{1}_{K_{j}}(u)~{}\mathrm{d}m_{K}(u)}{\int_{K}f_{j}(u)~{}\mathrm{d}m_{K}(u)}\bigg{(}1-\frac{1}{2^{j}}\bigg{)}.$
Observe that for all $u\in K$, $\mathbbm{1}_{\texttt{{B}}_{K}(w,\delta_{j})}(uw)=\mathbbm{1}_{\texttt{{B}}_{K}(e_{K},\delta_{j})}(u)$, so we have that
$\int_{K}\mathbbm{1}_{\texttt{{B}}_{K}(w,\delta_{j})}(u)\mathbbm{1}_{K_{j}}(u)~{}\mathrm{d}m_{K}(u)$ $=\int_{K}\mathbbm{1}_{\texttt{{B}}_{K}(w,\delta_{j})}(uw)\mathbbm{1}_{K_{j}}(uw)~{}\mathrm{d}m_{K}(u)$
$=\int_{K}\mathbbm{1}_{\texttt{{B}}_{K}(e_{K},\delta_{j})}(u)\mathbbm{1}_{K_{j}}(uw)~{}\mathrm{d}m_{K}(u)\geq\int_{K}f_{j}(u)\mathbbm{1}_{K_{j}}(uw)~{}\mathrm{d}m_{K}(u)$
$=\int_{K}f_{j}(uw^{-1})\mathbbm{1}_{K_{j}}(u)~{}\mathrm{d}m_{K}(u)=\int_{K_{j}}f_{j}(uw^{-1})~{}\mathrm{d}m_{K}(u).$
Combining the last equation with (4.15) and (4.16), we get that
$\frac{m_{K}(\{k\in K\colon\eta_{kH}(U)>0\}\cap\texttt{{B}}_{K}(w,\delta_{j}))}{m_{K}(\texttt{{B}}_{K}(w,\delta_{j}))}$ $\geq\frac{\int_{K_{j}}f_{j}(uw^{-1})~{}\mathrm{d}m_{K}(u)}{\int_{K}f_{j}(u)~{}\mathrm{d}m_{K}(u)}\bigg{(}1-\frac{1}{2^{j}}\bigg{)}$
$\geq\bigg{(}1-\frac{1}{j}\bigg{)}\bigg{(}1-\frac{1}{2^{j}}\bigg{)},$
where the least inequality is due to the fact that $w\in W_{j}$. Then, taking the limit as $j\to\infty$ we obtain that
$\lim_{j\to\infty}\frac{m_{K}(\{k\in K\colon\eta_{kH}(U)>0\}\cap\texttt{{B}}_{K}(w,\delta_{j}))}{m_{K}(\texttt{{B}}_{K}(w,\delta_{j}))}=1$
and this concludes the proof. ∎
We are now ready to prove Theorem 4.9.
Proof of Theorem 4.9.
Let $S=\{(x_{1},x_{2})\in X\times X\colon(x_{1},x_{2})\in\texttt{{supp}}(\lambda_{(x_{1},x_{2})})\}$. By Lemma 4.11, $S$ is a Borel subset of $X\times X$. Consider a sequence $\delta_{j}\to 0$ such that Proposition 4.13 is satisfied, and let $L\subset X$ be the set of $x\in X$ that satisfy (4.11). By Proposition 4.13, $\mu(L)=1$. Following the argument in [KMRR23b, Proposition 3.11] we will show that $\sigma(L\times L)=1$ and $L\times L\subset S$. Consequently, we will have that $\sigma(S)=1$, concluding the proof.
We start by showing that $\sigma(L\times L)=1$. We write $L\times L=(L\times X)\cap(X\times L)$ and so it is enough to show that both sets in this intersection have full measure $\sigma$. By Proposition 4.5 (i), we have that $\sigma(L\times X)=\pi_{1}\sigma(L)=\mu(L)=1$. By Proposition 4.5 (ii), the measure $\pi_{2}\sigma$ is absolutely continuous with respect to $\mu$, and since $\mu(L)=1$, it follows that $\sigma(X\times L)=\pi_{2}\sigma(L)=1$.
To conclude the proof, we show that $L\times L\subset S$. Let $(x_{1},x_{2})\in L\times L$. To show that $(x_{1},x_{2})\in S$, it is enough to show that for all open neighborhoods $U_{1},U_{2}$ of $x_{1},x_{2}$ we have $\lambda_{(x_{1},x_{2})}(U_{1}\times U_{2})>0$.
Let $U_{1},U_{2}$ be open neighborhoods of $x_{1},x_{2}$ respectively. By writing $\lambda_{(x_{1},x_{2})}=\int_{K}\eta_{kH}\times\eta_{kk_{1}^{-1}k_{2}H}~{}\mathrm{d}m_{K}(k)$, where $k_{1},k_{2}\in K$ are such that $\pi(x_{1})=k_{1}H$, $\pi(x_{2})=k_{2}H$, we see that it suffices to show that the set $W=W(k_{1},k_{2}):=\{k\in K:\eta_{kH}(U_{1})>0\text{ and }\eta_{kk_{1}^{-1}k_{2}H}(U_{2})>0\}$ has positive measure $m_{K}$, for some choice of the $k_{1},k_{2}$ as above. By Proposition 4.13, we can choose the elements $k_{1},k_{2}\in K$ such that
(4.17) $\frac{m_{K}(\{k\in K:\eta_{kH}(U_{1})>0\}\cap\texttt{{B}}_{K}(k_{1},\delta))}{m_{K}(\texttt{{B}}_{K}(k_{1},\delta))}\geq\frac{3}{4}$
and
(4.18) $\frac{m_{K}(\{k\in K:\eta_{kH}(U_{2})>0\}\cap\texttt{{B}}_{K}(k_{2},\delta))}{m_{K}(\texttt{{B}}_{K}(k_{2},\delta))}\geq\frac{3}{4},$
for some $\delta>0$. Now using (4.18) along with the bi-invariance of both $d_{K}$ and $m_{K}$, we have that
$\frac{m_{K}(\{k\in K:\eta_{kk_{1}^{-1}k_{2}H}(U_{2})>0\}\cap\texttt{{B}}_{K}(k_{1},\delta))}{m_{K}(\texttt{{B}}_{K}(k_{1},\delta))}$
$=\frac{m_{K}(\{k\in K:\eta_{kH}(U_{2})>0\}\cdot k_{2}^{-1}k_{1}\cap\texttt{{B}}_{K}(k_{2},\delta)\cdot k_{2}^{-1}k_{1})}{m_{K}(\texttt{{B}}_{K}(k_{2},\delta)\cdot k_{2}^{-1}k_{1})}$
$=\frac{m_{K}(\{k\in K:\eta_{kH}(U_{2})>0\}\cap\texttt{{B}}_{K}(k_{2},\delta))}{m_{K}(\texttt{{B}}_{K}(k_{2},\delta))}$
(4.19) $\geq\frac{3}{4}.$
Combining (4.17) and (4.2) yields $\frac{m_{K}(W)}{m_{K}(\texttt{{B}}_{K}(k_{1},\delta))}\geq\tfrac{1}{2}$. This implies that $m_{K}(W)>0$ and concludes the proof. ∎
It remains to show Theorem 4.10. To this end, we need the following lemma, which is the analog of [KMRR23b, Lemma 3.7] in our setting.
Lemma 4.14.
For $\sigma$-almost every $(x_{1},x_{2})\in X\times X$, we have $\lambda_{(a,x_{1})}=\lambda_{(x_{1},x_{2})}$.
Proof.
By the definition of $\sigma$ and the fact that $z\mapsto\eta_{z}$ is a disintegration, it follows that for $\sigma$-almost every $(x_{1},x_{2})$, we have $\pi(x_{1})=w\pi(a)$ and $\pi(x_{2})=w\pi(x_{1})$, for some $w\in K$. For any such $(x_{1},x_{2})$, using the right invariance of $m_{K}$, we have
$\lambda_{(x_{1},x_{2})}$ $=\int_{X\times X}\eta_{k\pi(x_{1})}\times\eta_{k\pi(x_{2})}~{}\mathrm{d}m_{K}(k)=\int_{X\times X}\eta_{kw\pi(a)}\times\eta_{kw\pi(x_{1})}~{}\mathrm{d}m_{K}(k)$
$=\int_{X\times X}\eta_{k\pi(a)}\times\eta_{k\pi(x_{1})}~{}\mathrm{d}m_{K}(kw^{-1})=\int_{X\times X}\eta_{k\pi(a)}\times\eta_{k\pi(x_{1})}~{}\mathrm{d}m_{K}(k)=\lambda_{(a,x_{1})}.$
This concludes the proof. ∎
We are now ready to prove Theorem 4.10.
Proof of Theorem 4.10.
Consider the measure $\nu_{a}:=\delta_{a}\times\mu$, where $\delta_{a}$ denotes the Dirac mass at $a$.
Claim. There exists some Følner sequence $\Psi$ such that
$\nu_{a}(\{(x_{0},x_{1})\in X\times X:(x_{0},x_{1})\in\texttt{{gen}}(\lambda_{(x_{0},x_{1})},\Psi)\})=1.$
From the definition of $\nu_{a}$, it is clear that for $\nu_{a}$-almost every $(x_{0},x_{1})\in X\times X$, $x_{0}=a$. Therefore, assuming the Claim we have that
$\nu_{a}(\{(a,x_{1})\in X\times X:(a,x_{1})\in\texttt{{gen}}(\lambda_{(a,x_{1})},\Psi)\})=1,$
which, by the definition of $\nu_{a}$ implies that
$\mu(\{x_{1}\in X:(a,x_{1})\in\texttt{{gen}}(\lambda_{(a,x_{1})},\Psi)\})=1.$
Then, by Proposition 4.5 (i), we have that $\pi_{1}\sigma=\mu$, and therefore
$\sigma(\{(x_{1},x_{2})\in X\times X:(a,x_{1})\in\texttt{{gen}}(\lambda_{(a,x_{1})},\Psi)\})=1.$
From Lemma 4.14, we know that $\lambda_{(a,x_{1})}=\lambda_{(x_{1},x_{2})}$ for $\sigma$-almost every $(x_{1},x_{2})$, and consequently,
$\sigma(\{(x_{1},x_{2})\in X\times X:(a,x_{1})\in\texttt{{gen}}(\lambda_{(x_{1},x_{2})},\Psi)\})=1,$
which was to be proved. Now, to finish the proof of Theorem 4.10, it only remains to prove the Claim.
Proof of Claim.
In this proof we follow the argument used in the proof of [KMRR22, Theorem 7.6]. Apply Lemma 2.8 for the ergodic decomposition $(x_{0},x_{1})\mapsto\lambda_{(x_{0},x_{1})}$, to obtain a Følner sequence $\Phi$ such that
(4.20) $(\mu\times\mu)(\{(x_{0},x_{1})\in X\times X:(x_{0},x_{1})\in\texttt{{gen}}(\lambda_{(x_{0},x_{1})},\Phi)\})=1.$
Consider the map $X\to M(X)$, $s\mapsto\nu_{s}=\delta_{s}\times\mu$, where $\delta_{s}$ denotes the Dirac mass at $s$. It is quite straightforward to see that $s\mapsto\nu_{s}$ is a continuous disintegration of $\mu\times\mu$, and moreover, satisfies
(4.21) $(T_{g}\times T_{g})\nu_{s}=\nu_{T_{g}s}$
for any $s\in X$. By (4.20) and the fact that $s\mapsto\nu_{s}$ is a disintegration of $\mu\times\mu$, it follows that for $\mu$-almost every $s\in X$, $\nu_{s}$-almost every $(x_{0},x_{1})\in X\times X$ is in $\texttt{{gen}}(\lambda_{(x_{0},x_{1})},\Phi)$. Fix $b\in\texttt{{supp}}(\mu)$. Then
(4.22) $\nu_{b}\text{-almost every }(x_{0},x_{1})\text{ is in }\texttt{{gen}}(\lambda_{(x_{0},x_{1})},\Phi).$
By Lemma 2.5, there exists some sequence $(g_{n})_{n\in\mathbb{N}}$ in $G$ such that $T_{g_{n}}a\to b$, and now by continuity of the disintegration $s\mapsto\nu_{s}$, combined with (4.21), it follows that
(4.23) $(T_{g_{n}}\times T_{g_{n}})\nu_{a}\to\nu_{b}.$
Now, let $\mathcal{F}=(F_{k})_{k\in\mathbb{N}}$ be a dense subset of $C(X\times X)$ and for $k,N\in\mathbb{N}$, consider the sets
$A_{k,N}=\bigg{\{}(x_{0},x_{1})\in X\times X:\max_{1\leq j\leq k}\bigg{|}\frac{1}{|\Phi_{N}|}\sum_{g\in\Phi_{N}}F_{j}((T_{g}\times T_{g})(x_{0},x_{1}))-\int_{X\times X}F_{j}~{}\mathrm{d}\lambda_{(x_{0},x_{1})}\bigg{|}\leq\frac{1}{k}\bigg{\}}.$
Using (4.22) and the monotone convergence theorem, it follows that for any $k\in\mathbb{N}$ there exists some $N(k)\in\mathbb{N}$, such that
(4.24) $\nu_{b}(A_{k,N(k)})\geq 1-2^{-k}.$
For $k,N\in\mathbb{N}$, we define
$B_{k,N}=\bigg{\{}(x_{0},x_{1})\in X\times X:d_{X\times X}((x_{0},x_{1}),A_{k,N})<\frac{1}{k}\bigg{\}},$
where $d_{X\times X}$ is the metric on $X\times X$. Then for $k$ sufficiently large, we have
(4.25) $\max_{1\leq j\leq k}\bigg{|}\frac{1}{|\Phi_{N(k)}|}\sum_{g\in\Phi_{N(k)}}F_{j}((T_{g}\times T_{g})(x_{0},x_{1}))-\int_{X\times X}F_{j}~{}\mathrm{d}\lambda_{(x_{0},x_{1})}\bigg{|}\leq\frac{2}{k},$
for any $(x_{0},x_{1})\in B_{k,N(k)}$. The sets $A_{k,N}$ are open, while the sets $B_{k,N}$ are closed subsets of $X\times X$, and also $A_{k,N(k)}\subset B_{k,N(k)}$, so by Urysohn’s lemma, we can find, for all $k\in\mathbb{N}$, continuous functions $f_{k}:X\times X\to[0,1]$ such that
$f_{k}|_{A_{k,N(k)}}=1\text{\quad and\quad}f_{k}|_{(X\times X)\setminus B_{k,N(k)}}=0.$
By (4.22), for each $k\in\mathbb{N}$, there exists $n(k)\in\mathbb{N}$ such that
$\bigg{|}\int_{X\times X}(T_{g_{n(k)}}\times T_{g_{n(k)}})f_{k}~{}\mathrm{d}\nu_{a}-\int_{X\times X}f_{k}~{}\mathrm{d}\nu_{b}\bigg{|}\leq 2^{-k}.$
Let $(h_{k})_{k\in\mathbb{N}}$ be the subsequence of $(g_{n})_{n\in\mathbb{N}}$ defined by $h_{k}=g_{n(k)}$, for $k\in\mathbb{N}$. Then, by the equation above, we have
$\nu_{a}\big{(}(T_{h_{k}}\times T_{h_{k}})^{-1}B_{k,N(k)}\big{)}$ $\geq\int_{X\times X}(T_{h_{k}}\times T_{h_{k}})f_{k}~{}\mathrm{d}\nu_{a}\geq\int_{X\times X}f_{k}~{}\mathrm{d}\nu_{b}-2^{-k}$
[…]
for any $k\in\mathbb{N}$. Therefore, it holds that
$\sum_{k\geq 1}\nu_{a}\big{(}(T_{h_{k}}\times T_{h_{k}})^{-1}B_{k,N(k)}\big{)}=\infty,$
[…]
$\max_{1\leq j\leq k}\bigg{|}\frac{1}{|\Psi_{k}|}\sum_{g\in\Psi_{k}}F_{j}((T_{g}\times T_{g})(x_{0},x_{1}))-\int_{X\times X}F_{j}~{}\mathrm{d}\lambda_{(T_{h_{k}}\times T_{h_{k}})(x_{0},x_{1})}\bigg{|}\leq\frac{2}{k},$
where $\Psi$ is the Følner sequence defined by $\Psi_{k}=h_{k}\Phi_{N(k)}$, for $k\in\mathbb{N}$. Using (4.6), the above equation becomes
$\max_{1\leq j\leq k}\bigg{|}\frac{1}{|\Psi_{k}|}\sum_{g\in\Psi_{k}}F_{j}((T_{g}\times T_{g})(x_{0},x_{1}))-\int_{X\times X}F_{j}~{}\mathrm{d}\lambda_{(x_{0},x_{1})}\bigg{|}\leq\frac{2}{k}.$
Sending $k\to\infty$, we have shown that
$\lim_{k\to\infty}\frac{1}{|\Psi_{k}|}\sum_{g\in\Psi_{k}}F((T_{g}\times T_{g})(x_{0},x_{1}))=\int_{X\times X}F~{}\mathrm{d}\lambda_{(x_{0},x_{1})}$
holds for $\nu_{a}$-almost every $(x_{0},x_{1})\in X\times X$ and for any $F\in\mathcal{F}$. An approximation argument concludes the proof of the Claim. ∎
The proof of the theorem is complete. ∎
4.3. The proof of Theorem 3.8
We are now ready to prove Theorem 3.8.
Let $G$ be a square absolutely continuous group, let $(X,\mu,T)$ be an ergodic $G$-system admitting a continuous factor map to its Kronecker factor, and let $a\in\texttt{{gen}}(\mu,\Phi)$ for some Følner sequence $\Phi$. Let also $E$ be a clopen subset of $X$ with $\mu(E)>0$. Consider the measure $\sigma$ given in (4.1). By Theorems 4.9 and 4.10, we have that there is a Følner sequence $\Psi$ such that
(4.26) $\sigma(\{(x_{1},x_{2}):(a,x_{1})\in\texttt{{gen}}(\lambda_{(x_{1},x_{2})},\Psi)\text{ and }(x_{1},x_{2})\in\texttt{{supp}}(\lambda_{(x_{1},x_{2})})\})=1.$
On the other hand, by Theorem 4.6 we have that
(4.27) $\sigma\bigg{(}E\times\bigcup_{t\in G}T_{t}^{-1}E\bigg{)}>0.$
Combining (4.26) and (4.27) we get that there exists $(x_{1},x_{2})\in X\times X$ such that for the $T\times T$-invariant measure $\lambda:=\lambda_{(x_{1},x_{2})}$ we have that $(a,x_{1})\in\texttt{{gen}}(\lambda,\Psi)$, $(x_{1},x_{2})\in\texttt{{supp}}(\lambda)$ and also $(x_{1},x_{2})\in E\times\big{(}\bigcup_{t\in G}T_{t}^{-1}E\big{)}$. Hence, there is $t\in G$ such that $x_{1}\in E$ and $x_{2}\in T_{t}^{-1}E$. Finally, applying Lemma 2.5 for $Y=X\times X$, $S=T\times T$, $y=(a,x_{1})$ and $w=(x_{1},x_{2})$, $\nu=\lambda$ and for the Følner sequence $\Psi$ we get that there is an infinite sequence $(g_{n})_{n\in N}$ in $G$ such that $(T_{g_{n}}\times T_{g_{n}})(a,x_{1})\to(x_{1},x_{2})$ in $X\times X$. Therefore, $(a,x_{1},x_{2})\in X^{3}$ forms an Erdős progression, and this concludes the proof of Theorem 3.8.
5. Proof of the corollaries of Theorem 1.5
We start by showing Theorem 1.7, and then we will prove Corollaries 1.8 and 1.9. We split the proof of Theorem 1.7 into the following three lemmas:
Lemma 5.1.
Let $G$ be an amenable group, let $M>0$ and let $\phi:G\to G$ be a map such that for every $g\in G$, $|\phi^{-1}(\{g\})|\leq M$. Suppose that there exist two Følner sequences $\Phi$ and $\Psi$ in $G$ and some $\eta>0$ such that for any $N\in\mathbb{N}$, we have that $\phi(\Psi_{N})\subset\Phi_{N}$ and
$\frac{|\phi(\Psi_{N})|}{|\Phi_{N}|}\geq\eta.$
Then $G$ is $\phi$-absolutely continuous.
Lemma 5.2.
Let $G$ be a torsion-free finitely generated nilpotent group. Then the squaring map $s_{G}$ on $G$ is injective.
Lemma 5.3.
Let $G$ be a torsion-free finitely generated nilpotent group. Then there exists some Følner sequence $\Psi$ in $G$ and some $\eta>0$ such that for any $N\in\mathbb{N}$, we have that $s_{G}(\Psi_{N})\subset\Psi_{N+1}$ and
(5.1) $\frac{|s_{G}(\Psi_{N})|}{|\Psi_{N+1}|}=\eta.$
It is clear that Theorem 1.7 follows immediately from Lemmas 5.1, 5.2 and 5.3, where Lemma 5.1 is applied for $\phi=s_{G}$, $\Psi$ the Følner guaranteed by Lemma 5.3 and $\Phi$ the Følner given by $\Phi_{N}=\Psi_{N+1}$.
Proof of Lemma 5.1.
Let $G,\phi,\Phi,\Psi$ and $\eta$ be as in the assumptions of Lemma 5.1. We will prove something stronger than we require, namely that for any $u:G\to[0,1]$ we have
$\limsup_{N\to\infty}\frac{1}{|\Psi_{N}|}\sum_{g\in\Psi_{N}}u(\phi(g))\leq\frac{M}{\eta}\limsup_{N\to\infty}\frac{1}{|\Phi_{N}|}\sum_{g\in\Phi_{N}}u(g),$
and then for any $\varepsilon>0$, taking $\delta=\eta\varepsilon/M>0$ yields that $G$ is $\phi$-absolutely continuous. Let $u:G\to[0,1]$, and let $(\Psi_{N_{k}})_{k\in\mathbb{N}}$ such that
$\lim_{k\to\infty}\frac{1}{|\Psi_{N_{k}}|}\sum_{g\in\Psi_{N_{k}}}u(\phi(g))=\limsup_{N\to\infty}\frac{1}{|\Psi_{N}|}\sum_{g\in\Psi_{N}}u(\phi(g)).$
For each $k\in\mathbb{N}$, using the assumption on $\Phi$ and $\Psi$, we have that $\phi(\Psi_{N_{k}})\subset\Phi_{N_{k}}$ and $\frac{|\phi(\Psi_{N_{k}})|}{|\Phi_{N_{k}}|}\geq\eta$. Then for every $k\in\mathbb{N}$ we have
$\frac{1}{|\Psi_{N_{k}}|}\sum_{g\in\Psi_{N_{k}}}u(\phi(g))\leq\frac{M}{|\phi(\Psi_{N_{k}})|}\sum_{g\in\phi(\Psi_{N_{k}})}u(g)\leq\frac{M}{\eta}\frac{1}{|\Phi_{N_{k}}|}\sum_{g\in\Phi_{N_{k}}}u(g),$
so letting $k\to\infty$ we obtain that
$\lim_{k\to\infty}\frac{1}{|\Psi_{N_{k}}|}\sum_{g\in\Psi_{N_{k}}}u(\phi(g))\leq\frac{M}{\eta}\limsup_{k\to\infty}\frac{1}{|\Phi_{N_{k}}|}\sum_{g\in\Phi_{N_{k}}}u(g)\leq\frac{M}{\eta}\limsup_{N\to\infty}\frac{1}{|\Phi_{N}|}\sum_{g\in\Phi_{N}}u(g),$
which concludes the proof. ∎
Let $G$ be a torsion-free finitely generated nilpotent group and fix a Mal’cev coordinate system $(t_{1},\dots,t_{s})$. By [KM79, Theorem 17.2.5], we have that for any $1\leq i\leq s$, there are polynomials $p_{i}$ in $i-1$ variables, with rational coefficients, satisfying $p_{i}(\mathbb{Z}^{i-1})\subset\mathbb{Z}$ such that for any $x\in G$, we have
(5.2) $t_{i}(x^{2})=2t_{i}(x)+p_{i}(\{t_{j}(x)\colon j<i\}).$
Proof of Lemma 5.2.
The proof becomes obvious by using (5.2). ∎
Proof of Lemma 5.3.
Let $G$ be a torsion-free finitely generated nilpotent group and fix a Mal’cev coordinate system $(t_{1},\dots,t_{s})$. We identify the group $G$ with $\mathbb{Z}^{s}$ as implied by the Mal’cev coordinate system. Let $(p_{i})_{1\leq i\leq s}$ be the sequence of polynomials satisfying (5.2) and for each $i$ we denote the number of terms of $p_{i}$ by $\gamma_{i}$. Now we will define three integer-valued sequences $(b_{i})_{1\leq i\leq s},(c_{i})_{1\leq i\leq s}$ and $(d_{i})_{1\leq i\leq s}$ that will help us to construct the Følner sequence $\Psi$ with the required properties. We start with the latter one and we define it recursively as follows: Let $d_{1}=1$. Now let $1<i\leq s$ and suppose that $d_{j}$ has been defined for all $1\leq j<i$. Given a monomial $m(x_{1},\dots,x_{i-1})=x_{1}^{e_{1}}\dots x_{i-1}^{e_{i-1}}$, with $e_{j}\geq 0$ for any $1\leq j<i$, we let $d(m)=\sum_{1\leq j<i}e_{j}d_{j}$, and then we also let $d(p_{i})=\max\{d(m)\colon\text{the monomial}~{}m~{}\text{appears in}~{}p_{i}\}$. Then we define $d_{i}=\max\{d(p_{i}),1\}$. The other two sequences are defined as follows: Let $b_{1}=0$. Now we fix $1<i\leq s$. We denote by $m_{i}$ the monomial $m$ appearing in $p_{i}$ with maximal coefficient in absolute value such that $d(p_{i})=d(m)$. Then we define $b_{i}$ to be ceiling of the absolute value of the coefficient of $m_{i}$. Finally, for any $1\leq i\leq s$, we let $c_{i}=\gamma_{i}b_{i}+2$. We also note that $d_{i}=d(m_{i})$ for each $1\leq i\leq s$.
Let us now define the Følner sequence $\Psi$. By the definition of $(b_{i})_{1\leq i\leq s}$ and $(d_{i})_{1\leq i\leq s}$, we have that for any $1\leq i\leq s$, if $|x_{j}|\leq M^{d_{j}}$ for every $1\leq j<i$, for some $M>0$, then $|m_{i}(x_{1},\dots,x_{i-1})|\leq b_{i}M^{d_{i}}$, and then, $|p_{i}(x_{1},\dots,x_{i-1})|\leq\gamma_{i}b_{i}M^{d_{i}}$. It follows that for any $1\leq i\leq s$, the following implication holds:
[…]
Given $M>0$, we use the notation $[M]:=(-M,M]\cap\mathbb{Z}$. Now, for any $N\in\mathbb{N}$, we define
$\Psi_{N}=[c_{1}^{(N-1)d_{1}}]\times[c_{2}^{(N-1)d_{2}}]\times\dots\times[c_{s}^{(N-1)d_{s}}].$
It is not hard to check that $\Psi=(\Psi_{N})_{N\in\mathbb{N}}$ is Følner sequence in $G$. We show that $s(\Psi_{N})\subset\Psi_{N+1}$ for every $N\in\mathbb{N}$. Let $N\in\mathbb{N}$ and then
[…]
Let $x^{2}=(t_{1}(x^{2}),\dots,t_{s}(x^{2}))\in s(\Psi_{N})$. We claim that for each $1\leq i\leq s$, we have that
(5.4) $t_{i}(x^{2})\in[c_{i}^{Nd_{i}}].$
Let $1\leq i\leq s$. For any $1\leq j<i$, we have that $t_{j}(x)\in[c_{j}^{(N-1)d_{j}}]$, and then we have that
$t_{i}(x^{2})=2t_{i}(x)+p_{i}((t_{j}(x)\colon j<i))$
$\in(p_{i}((t_{j}(x))_{1\leq j<i})-2c_{i}^{(N-1)d_{i}},p_{i}((t_{j}(x))_{1\leq j<i})+2c_{i}^{(N-1)d_{i}}]\cap(2\mathbb{Z}+p_{i}((t_{j}(x))_{1\leq j<i}))$
$\subset[c_{i}^{Nd_{i}}],$
where the last one follows from that $|p_{i}((t_{j}(x))_{1\leq j<i})|\leq\gamma_{i}b_{i}c_{i}^{(N-1)d_{i}}$, by (5.3), and from the definition of $c_{i}$ along with that $d_{i}\geq 1$. This shows (5.4) and hence that $s(\Psi_{N})\subset\Psi_{N+1}$.
It remains to prove (5.1) and we show it for the Følner $\Psi$ and with $\eta:=\prod_{1\leq i\leq s}c_{i}^{-d_{i}}>0$. In other words, we prove that for any $N\in\mathbb{N}$ we have
(5.5) $\frac{|s(\Psi_{N})|}{|\Psi_{N+1}|}=\prod_{1\leq i\leq s}c_{i}^{-d_{i}}.$
Let $N\in\mathbb{N}$. By Lemma 5.2, $s_{G}$ is injective, and since $\Psi_{N}$ is finite, we have that
$|s(\Psi_{N})|=|\Psi_{N}|=\prod_{1\leq i\leq s}2^{s}\bigg{(}\prod_{1\leq i\leq s}c_{i}^{d_{i}}\bigg{)}^{N-1}.$
Moreover, we clearly have
$|\Psi_{N+1}|=\prod_{1\leq i\leq s}2^{s}\bigg{(}\prod_{1\leq i\leq s}c_{i}^{d_{i}}\bigg{)}^{N},$
and so (5.5) follows from comparing the last two equations. This concludes the proof. ∎
Having established Theorem 1.7, we are ready to prove Corollary 1.8. Before we prove it, let us make some remarks.
Let $s\in\mathbb{N}$ and $1\leq i\leq s$. We say that a non-empty set $L\subset\mathbb{Z}^{s}$ is a line in the $i$-th coordinate if there is a non-empty interval $I$ in $\mathbb{Z}$, i.e. a set of the form $\{m,\dots,m+r\}$ for some $m,r\in\mathbb{Z}$ with $r\geq 0$, and some integers $x_{j}$, $j\in\{1,\dots,s\}\setminus\{i\}$ such that
$L=\{x_{1}\}\times\dots\times\{x_{i-1}\}\times I\times\{x_{i+1}\}\times\dots\times\{x_{s}\}.$
We refer to the length of the interval $I$ as the length of the line $L$.
Let $G$ be a torsion-free finitely generated nilpotent group and $(t_{1},\dots,t_{s})$ a Mal’cev coordinate system. For each $1\leq i\leq s$, let $e_{i}$ be the element with coordinates $(e^{(i)}_{1},\dots,e^{(i)}_{s})$, where $e^{(i)}_{i}=1$ and $e^{(i)}_{j}=0$ for $j\neq i$. Every $g\in G$ has a unique representation as $(t_{1}(g),\dots,t_{s}(g))\in\mathbb{Z}^{s}$, and this defines a bijective map from $G$ to $\mathbb{Z}^{s}$. From now on, we identify each $g\in G$ with its coordinates $(t_{1}(g),\dots,t_{s}(g))\in\mathbb{Z}^{s}$, and every set $E\subset G$ with the corresponding set of coordinates in $\mathbb{Z}^{s}$. We freely pass from viewing a set $E\subset G$ as a subset of $\mathbb{Z}^{s}$ and vice versa, without stating it, as it will be clear from the context.
If $E$ is a finite subset of $G$ and $1\leq i\leq s$, then $E$ can be written as a finite disjoint union of lines in the $i$-th coordinate, and this can be done in many ways. We want to write $E$ as a union of lines which is going to be maximal in some sense that is going to be useful for us in our proof of Corollary 1.8.
More precisely, if $E$ is a finite subset of $G$, and $1\leq i\leq s$, then we can always write it as a disjoint union $E=\bigsqcup_{j=1}^{\ell}L^{(j)}$ such that the sets $L^{(j)}$ are lines in the $i$-th coordinate, and for each $j$, the line $L^{(j)}$ is maximal within $E$, meaning that for each $1\leq j\leq s$, $e_{i}L^{(j)}$ is not a subset of $E$. Note that although there always exists such a choice of lines, it may not be unique, but uniqueness is not necessary for our purposes. We are now ready to prove Corollary 1.8.
Proof of Corollary 1.8.
The first part of Corollary 1.8 follows immediately by combining Theorems 1.5 and 1.7. It remains to prove that if $G$ is a torsion-free finitely generated nilpotent group, then given a Mal’cev coordinate system $(t_{1},\dots,t_{s})$ on $G$, we can choose $B$ so that for any finite set $C\subset\mathbb{Z}$ and any $1\leq i\leq s$, the set $\{b\in B:t_{i}(b)\in C\}$ is finite.
Let $G$ be a torsion-free finitely generated nilpotent group, let $A$ have positive left upper Banach density, and $\Psi=(\Psi_{N})_{N\in\mathbb{N}}$ a left Følner sequence such that $d_{\Psi}(A)>0$, where the previous density exists. In addition, let $(t_{1},\dots,t_{s})$ be any Mal’cev coordinate system on $G$.
Claim. There is a Følner sequence $\Psi^{\prime}$ such that $d_{\Psi^{\prime}}(A)>0$ and for any finite set $C\subset\mathbb{Z}$ and any $1\leq i\leq s$, the set $\{N\in\mathbb{N}:t_{i}(\Psi^{\prime}_{N})\cap C\neq\emptyset\}$ is finite.
[…]
Proof of Claim.
We know that $d_{\Psi}(A)=\lim_{N\to\infty}\frac{|A\cap\Psi_{N}|}{|\Psi_{N}|}>0$. For each $1\leq i\leq s$ and $N\in\mathbb{N}$, let
$\delta_{i,N}=\frac{|\Psi_{N}\triangle(e_{i}\Psi_{N})|}{|\Psi_{N}|}.$
Since $\Psi$ is a Følner sequence, for all $i$, $\delta_{i,N}\to 0$ as $N\to\infty$. Then we can choose a sequence $Q_{N}$ of natural numbers such that $Q_{N}\to\infty$ as $N\to\infty$ and for all $i$, $Q_{N}\delta_{i,N}\to 0$ as $N\to\infty$.
Step $1$: For each $N\in\mathbb{N}$, $\Psi_{N}$ is a finite subset of $G$, so we can write it as a disjoint union of lines in the $1$st coordinate, which are maximal within $\Psi_{N}$. Let $\Psi_{N}^{(1)}$ be the union of those lines whose length is greater that $Q_{N}$, and $m_{N}$ be the number of those lines whose length is less than or equal to $Q_{N}$. Then
$Q_{N}\delta_{1,N}=Q_{N}\frac{|\Psi_{N}\triangle(e_{1}\Psi_{N})|}{|\Psi_{N}|}\geq\frac{Q_{N}m_{N}}{|\Psi_{N}|}\geq\frac{|\Psi_{N}|-|\Psi_{N}^{(1)}|}{|\Psi_{N}|}.$
Therefore, $\frac{|\Psi_{N}^{(1)}|}{|\Psi_{N}|}\to 1$ as $N\to\infty$, from which one gets that $\Psi^{(1)}=(\Psi_{N}^{(1)})_{N\in\mathbb{N}}$ is a left Følner sequence in $G$. In addition, it is not difficult to see that the density $d_{\Psi^{(1)}}(A)$ exists and $d_{\Psi^{(1)}}(A)=d_{\Psi}(A)$.
Recall that for each $N\in\mathbb{N}$, $\Psi_{N}^{(1)}$ is a disjoint union of some lines in the $1$-st coordinate $L^{(1,N)},\dots,L^{(\ell_{N},N)}$ whose length is greater than $Q_{N}$. Let $N_{1}=1$ and set $\widetilde{\Psi}_{1}^{(1)}=\Psi_{1}^{(1)}$. Since $\Psi_{1}^{(1)}$ is finite, the projection $P_{1}$ of $\Psi_{1}^{(1)}$ in the first coordinate is also finite. For each $N\in\mathbb{N}$, let
$\Psi_{N}^{(1,2)}=\{g\in\Psi_{N}^{(1)}:t_{1}(g)\notin P_{1}\}=\bigsqcup_{j=1}^{\ell_{N}}\{g\in L^{(j,N)}:t_{1}(g)\notin P_{1}\}.$
Then $\Psi_{N}^{(1,2)}\subset\Psi_{N}^{(1)}$ and
$\frac{|\Psi_{N}^{(1,2)}|}{|\Psi_{N}^{(1)}|}$ $=\frac{\sum_{j=1}^{\ell_{N}}|\{g\in L^{(j,N)}:t_{1}(g)\notin P_{1}\}|}{\sum_{j=1}^{\ell_{N}}|L^{(j,N)}|}\geq\frac{\sum_{j=1}^{\ell_{N}}(|L^{(j,N)}|-|P_{1}|)}{\sum_{j=1}^{\ell_{N}}|L^{(j,N)}|}$
$=1-\frac{\sum_{j=1}^{\ell_{N}}|P_{1}|}{\sum_{j=1}^{\ell_{N}}|L^{(j,N)}|}\geq 1-\frac{\ell_{N}|P_{1}|}{\ell_{N}Q_{N}}=1-\frac{|P_{1}|}{Q_{N}}.$
Since $Q_{N}\to\infty$ as $N\to\infty$, we have that $\frac{|P_{1}|}{Q_{N}}\to 0$ as $N\to\infty$, so $\frac{|\Psi_{N}^{(1,2)}|}{|\Psi_{N}^{(1)}|}\to 1$ as $N\to\infty$. Hence, we can pick $N_{2}\in N,N_{2}>N_{1}$ such that $\frac{|\Psi_{N_{2}}^{(1,2)}|}{|\Psi_{N_{2}}^{(1)}|}>\frac{1}{2}$. Set $\widetilde{\Psi}^{(1)}_{2}=\Psi_{N_{2}}^{(1,2)}.$
Now since $\widetilde{\Psi}^{(1)}_{2}$ is finite, the projection $P_{2}$ of $\widetilde{\Psi}_{2}^{(1)}$ in the first coordinate is also finite. For each $N\in\mathbb{N}$, let
$\Psi_{N}^{(1,3)}=\{g\in\Psi_{N}^{(1)}:t_{1}(g)\notin P_{1}\cup P_{2}\}=\bigsqcup_{j=1}^{\ell_{N}}\{g\in L^{(j,N)}:t_{1}(g)\notin P_{1}\cup P_{2}\}.$
Then $\Psi_{N}^{(1,3)}\subset\Psi_{N}^{(1)}$ and as before, we have that
$\frac{|\Psi_{N}^{(1,3)}|}{|\Psi_{N}^{(1)}|}\geq 1-\frac{|P_{1}|+|P_{2}|}{Q_{N}}.$
Again, since $Q_{N}\to\infty$ as $N\to\infty$, we have that $\frac{|P_{1}|+|P_{2}|}{Q_{N}}\to 0$ as $N\to\infty$, so $\frac{|\Psi_{N}^{(1,3)}|}{|\Psi_{N}^{(1)}|}\to 1$ as $N\to\infty$. Hence, we can pick $N_{3}\in N,N_{3}>N_{2}$ such that $\frac{|\Psi_{N_{3}}^{(1,3)}|}{|\Psi_{N_{3}}^{(1)}|}>\frac{2}{3}$. Set $\widetilde{\Psi}^{(1)}_{3}=\Psi_{N_{3}}^{(1,3)}.$
Continuing inductively, we find a strictly increasing sequence of natural numbers $(N_{k})_{k\in\mathbb{N}}$ such that for all $k\in\mathbb{N}$, $\widetilde{\Psi}^{(1)}_{k}\subset{\Psi}^{(1)}_{N_{k}}\subset{\Psi}_{N_{k}}$ and $\frac{|\widetilde{\Psi}^{(1)}_{k}|}{|\Psi_{N_{k}}^{(1)}|}>1-\frac{1}{k}$. Then $\frac{|\widetilde{\Psi}^{(1)}_{k}|}{|\Psi_{N_{k}}^{(1)}|}\to 1$ as $k\to\infty$, from which one gets that $\widetilde{\Psi}^{(1)}=(\widetilde{\Psi}_{k}^{(1)})_{k\in\mathbb{N}}$ is a left Følner sequence in $G$. In addition, it is not difficult to see that the density $d_{\widetilde{\Psi}^{(1)}}(A)$ exists and $d_{\widetilde{\Psi}^{(1)}}(A)=d_{({\Psi}^{(1)}_{N_{k}})_{k\in\mathbb{N}}}(A)=d_{\Psi^{(1)}}(A)=d_{\Psi}(A)$. In addition, from the construction of $\widetilde{\Psi}^{(1)}$, we have that for any finite set $C\subset\mathbb{Z}$, the set $\{k\in\mathbb{N}:t_{1}(\widetilde{\Psi}^{(1)}_{k})\cap C\neq\emptyset\}$ is finite.
Step $2$: Repeat Step 1 with $\widetilde{\Psi}^{(1)}$ in place of $\Psi$, which we write as a disjoint union of lines in the $2$nd coordinate that are maximal within $\widetilde{\Psi}_{N}^{(1)}$, to obtain a strictly increasing sequence of natural numbers $(N_{k})_{k\in\mathbb{N}}$ and a left Følner sequence $\widetilde{\Psi}^{(2)}$ such that for all $k\in\mathbb{N}$, $\widetilde{\Psi}^{(2)}_{k}\subset\widetilde{\Psi}^{(1)}_{N_{k}}$, $\frac{|\widetilde{\Psi}^{(2)}_{k}|}{|\widetilde{\Psi}^{(1)}_{N_{k}}|}\to 1$ as $k\to\infty$ and such that for any finite set $C\subset\mathbb{Z}$, the set $\{k\in\mathbb{N}:t_{2}(\widetilde{\Psi}^{(2)}_{k})\cap C\neq\emptyset\}$ is finite. Then we will also have that the density $d_{\widetilde{\Psi}^{(2)}}(A)$ exists and $d_{\widetilde{\Psi}^{(2)}}(A)=d_{\widetilde{\Psi}^{(1)}}(A)=d_{\Psi}(A)$.
Recall that $\widetilde{\Psi}^{(1)}$ has the property that for any finite set $C\subset\mathbb{Z}$, the set $\{N\in\mathbb{N}:t_{1}(\widetilde{\Psi}^{(1)}_{N})\cap C\neq\emptyset\}$ is finite. As $\widetilde{\Psi}^{(2)}_{k}\subset\widetilde{\Psi}^{(1)}_{N_{k}}$ for all $k\in\mathbb{N}$, we get that for any finite set $C\subset\mathbb{Z}$, the set $\{k\in\mathbb{N}:t_{1}(\widetilde{\Psi}^{(2)}_{k})\cap C\neq\emptyset\}$ is finite. Hence, after all, for any finite set $C\subset\mathbb{Z}$ and any $i\in\{1,2\}$, the set $\{k\in\mathbb{N}:t_{i}(\widetilde{\Psi}^{(2)}_{k})\cap C\neq\emptyset\}$ is finite.
Repeating the same procedure, after $s$ steps, we find a left Følner sequence $\widetilde{\Psi}^{(s)}$ such that the density $d_{\widetilde{\Psi}^{(s)}}(A)$ exists, $d_{\widetilde{\Psi}^{(s)}}(A)=d_{\Psi}(A)>0$ and for any finite set $C\subset\mathbb{Z}$ and any $i\in\{1,\dots,s\}$, the set $\{N\in\mathbb{N}:t_{i}(\widetilde{\Psi}^{(s)}_{N})\cap C\neq\emptyset\}$ is finite. Taking $\Psi^{\prime}:=\widetilde{\Psi}^{(s)}$, we see that $\Psi^{\prime}$ satisfies the Claim, thus concludes its proof. ∎
Since the Claim is established, the proof of the corollary is complete. ∎
Now, we have to show Corollary 1.9, but this is not hard using the fact that finitely generated nilpotent groups are virtually torsion-free. Let us first show the following simple lemma:
Lemma 5.4.
Let $G$ be an amenable group, $H$ be a subgroup of $G$ with $|G:H|=r<\infty$ and $\Phi$ be a Følner sequence in $G$. Then the following hold:
  • (i)
    $d_{\Phi}(H)=\frac{1}{r}$.
  • (ii)
    Letting $\Psi_{N}=\Phi_{N}\cap H$ for each $N\in\mathbb{N}$, the sequence $\Psi=(\Psi_{N})_{N\in\mathbb{N}}$ is a Følner sequence in $H$.
In the following proof and onwards, we use the symbol $\sqcup$ to denote the disjoint union.
Proof.
(i) This is quite easy to check.
(ii) Let $h\in H$ and suppose for sake of contradiction that
$\ell=\liminf_{N\to\infty}\frac{|h\Psi_{N}\cap\Psi_{N}|}{|\Psi_{N}|}<1.$
For any $N\in\mathbb{N}$, we have that
$h\Phi_{N}\cap\Phi_{N}=(h\Psi_{N}\cap\Psi_{N})\sqcup((h\Phi_{N}\cap\Phi_{N})\setminus H),$
hence,
$\frac{|h\Phi_{N}\cap\Phi_{N}|}{|\Phi_{N}|}=\frac{|h\Psi_{N}\cap\Psi_{N}|}{|\Phi_{N}|}+\frac{|(h\Phi_{N}\cap\Phi_{N})\setminus H|}{|\Phi_{N}|}\leq\frac{|h\Psi_{N}\cap\Psi_{N}|}{|\Psi_{N}|}\cdot\frac{|\Psi_{N}|}{|\Phi_{N}|}+\frac{|\Phi_{N}\setminus H|}{|\Phi_{N}|}.$
Using (i), it follows that
$\lim_{N\to\infty}\frac{|h\Phi_{N}\cap\Phi_{N}|}{|\Phi_{N}|}\leq\frac{\ell}{r}+1-\frac{1}{r}<1,$
which contradicts the fact that $\Phi$ is a Følner sequence in $G$. Therefore, $\Psi$ is indeed a Følner sequence in $H$ and the proof of the lemma is complete. ∎
Proof of Corollary 1.9.
Let $G$ be a finitely generated virtually nilpotent group, $\Phi=(\Phi_{N})_{N\in\mathbb{N}}$ be a Følner sequence in $G$ and $A$ be a subset of $G$ with $\overline{d}_{\Phi}(A)>0$. By passing to a subsequence for which the limit exists, we may assume that $d_{\Phi}(A)>0$. Let $G^{\prime}$ be a nilpotent finite-index subgroup of $G$. Since $G$ is finitely generated, by writing it as a disjoint union of finitely many left cosets of $G^{\prime}$, it is easy to see that $G^{\prime}$ is also finitely generated, and it is also nilpotent. By [KM79, Theorem 17.2.2], there exists a normal subgroup $H$ of $G^{\prime}$ with finite index, which is torsion-free. Hence, $H$ is a torsion-free finitely generated nilpotent group, which has finite index in $G$. By writing $G$ as finite disjoint union of left cosets of $H$, we can see that there exists some $g\in G$ such that $\overline{d}_{\Phi}(g^{-1}A\cap H)=\overline{d}_{\Phi}(A\cap gH)>0$. Again, by passing to a subsequence, we may assume that the limit exists, so $d_{\Phi}(g^{-1}A\cap H)=d_{\Phi}(A\cap gH)>0$. We let $\Psi_{N}=\Phi_{N}\cap H$ for each $N\in\mathbb{N}$ and by Lemma 5.4 (ii), $\Psi=(\Psi_{N})_{N\in\mathbb{N}}$ is a Følner sequence on $H$. Hence for every $N\in\mathbb{N}$ we have that
$\frac{|g^{-1}A\cap\Psi_{N}|}{|\Psi_{N}|}=\frac{|g^{-1}A\cap\Phi_{N}\cap H|}{|\Phi_{N}|}\cdot\frac{|\Phi_{N}|}{|\Psi_{N}|},$
thus,
$d_{\Psi}(g^{-1}A\cap H)=\frac{d_{\Phi}(g^{-1}A\cap H)}{d_{\Phi}(H)}>0,$
[…]
Once we have established the corollaries concerning finitely generated nilpotent groups, we move on to showing the results corresponding to abelian groups, namely Theorem 1.11 and Corollary 1.12. Obviously, the latter is an immediate consequence of the former and of Theorem 1.5, so it suffices to show Theorem 1.11.
Proof of Theorem 1.11.
Let $(G,+)$ be an abelian group such that $2G$ has finite index in $G$, let $\Psi$ be any Følner sequence in $G$ and consider the doubling map $s_{G}:G\to 2G,g\mapsto 2g$. We will show that the assumptions of Lemma 5.1 are satisfied, which will yield that $G$ is square absolutely continuous. Then $|\ker(s_{G})|=|G:2G|=r$, hence
(5.6) $|s_{G}^{-1}(\{g\})|\leq r,\quad\forall g\in G.$
By assumption, there exist $r\in\mathbb{N}$ and $g_{1},\dots,g_{r}\in G$, where $g_{1}=e_{G}$, such that
$G=\bigsqcup_{i=1}^{r}(g_{i}+2G).$
We define $\Phi=(\Phi_{N})_{N\in\mathbb{N}}$ by $\Phi_{N}=\bigsqcup_{i=1}^{r}(g_{i}+2\Psi_{N})$. Assuming that this is a Følner sequence in $G$, by taking $\eta:=\frac{1}{r}$, we have that
(5.7) $s_{G}(\Psi_{N})=2\Psi_{N}\subset\Phi_{N}\qquad\text{and}\qquad\frac{|s_{G}(\Psi_{N})|}{|\Phi_{N}|}=\frac{|2\Psi_{N}|}{r|2\Psi_{N}|}=\eta.$
Combining (5.6) and (5.7), we have that the assumptions of Lemma 5.1 are indeed satisfied. Thus, it suffices to show that $\Phi$ is a Følner sequence in $G$.
Let $g\in G$ and $\varepsilon>0$. There exist $1\leq i_{0}\leq r$ and $h\in G$ such that $g=g_{i_{0}}+2h$. Then
(5.8) $(g+\Phi_{N})\cap\Phi_{N}=\bigg{(}\bigsqcup_{i=1}^{r}(g_{i_{0}}+g_{i}+2(h+\Psi_{N}))\bigg{)}\cap\bigg{(}\bigsqcup_{j=1}^{r}(g_{j}+2\Psi_{N})\bigg{)}.$
Since the cosets $g_{j}+2G$ are disjoint, it follows that for each $1\leq i\leq r$, there exists a unique $1\leq j(i)\leq r$ such that $g_{i_{0}}+g_{i}\in g_{j(i)}+2G$, so there is $y_{i}\in G$ such that $g_{i_{0}}+g_{i}=g_{j(i)}+2y_{i}$. In addition, if we assume that for $i_{1}\neq i_{2}$ we have $j(i_{1})=j(i_{2})$, then we have that $g_{i_{1}}-g_{i_{2}}=2y_{i_{1}}-2y_{i_{2}}\in 2G$, so $g_{i_{1}}+2G=g_{i_{2}}+2G$, which is a contradiction. Therefore, the map $j:\{1,\dots,r\}\to\{1,\dots,r\},i\mapsto j(i)$ is a bijection, and then (5.8) becomes
$(g+\Phi_{N})\cap\Phi_{N}=\bigsqcup_{i=1}^{r}\big{(}(g_{j(i)}+2(y_{i}+h+\Psi_{N}))\cap(g_{j(i)}+2\Psi_{N})\big{)}=\bigsqcup_{i=1}^{r}\big{(}g_{j(i)}+(2(y_{i}+h+\Psi_{N})\cap 2\Psi_{N})),$
thus, we have that
$|(g+\Phi_{N})\cap\Phi_{N}|$ $=\sum_{i=1}^{r}|2(y_{i}+h+\Psi_{N})\cap 2\Psi_{N})|=\sum_{i=1}^{r}|s_{G}(y_{i}+h+\Psi_{N})\cap s_{G}(\Psi_{N}))|$
(5.9) $\geq\sum_{i=1}^{r}|s_{G}((y_{i}+h+\Psi_{N})\cap\Psi_{N})|.$
Now since $\Psi$ is a Følner sequence in $G$, for $N$ sufficiently large, we have that for every $1\leq i\leq r$, $|(y_{i}+h+\Psi_{N})\triangle\Psi_{N}|\leq\frac{\varepsilon}{r}|\Psi_{N}|$, and then we have that
(5.10) $\frac{|s_{G}((y_{i}+h+\Psi_{N})\cap\Psi_{N})|}{|s_{G}(\Psi_{N})|}\geq 1-\frac{|s_{G}(\Psi_{N}\triangle(y_{i}+h+\Psi_{N}))|}{|s_{G}(\Psi_{N})|}\geq 1-\frac{r|\Psi_{N}\triangle(y_{i}+h+\Psi_{N})|}{|\Psi_{N}|}\geq 1-\varepsilon,$
where for the second inequality above we use that $|s_{G}(\Psi_{N}\triangle(y_{j}+h+\Psi_{N}))|\leq|(\Psi_{N}\triangle(y_{j}+h+\Psi_{N}))|$ and that $|s_{G}(\Psi_{N})|\geq\frac{1}{r}|\Psi_{N}|$, as $|s_{G}^{-1}(\{z\})|\leq r$ for all $z\in G$.
Then, combining (5) and (5.10) we get that for $N$ sufficiently large
$\frac{|(g+\Phi_{N})\cap\Phi_{N}|}{|\Phi_{N}|}\geq\frac{\sum_{i=1}^{r}(1-\varepsilon)|s_{G}(\Psi_{N})|}{r|s_{G}(\Psi_{N})|}=1-\varepsilon.$
Since $\varepsilon>0$ was arbitrary, it follows that $\lim_{N\to\infty}\frac{|(g+\Phi_{N})\cap\Phi_{N}|}{|\Phi_{N}|}=1$. Thus, $\Phi$ is a Følner sequence in $G$, and the proof is complete. ∎
6. Counterexamples on products sets
To construct the counterexamples we introduce some convenient notation. We denote by $H_{3}$ the $3\times 3$ discrete Heisenberg group, that is the group of $3\times 3$ upper triangular matrices with $1$ in the diagonal and integer entries. Using the obvious Mal’cev coordinate system in this group we identify $H_{3}$ with $\mathbb{Z}^{3}$ by writing elements of $H_{3}$ as $a=(a_{1},a_{2},a_{3})$ where the group operation is given by $ab=(a_{1}+b_{1},a_{2}+b_{2},a_{3}+b_{3}+a_{1}b_{2})$. For $N\in\mathbb{N}$, we denote $[N]:=[1,N]$, where all intervals are considered in $\mathbb{Z}$, and $[N]^{\prime}:=[1,N]\cap(2\mathbb{Z}+1)$. All the counterexamples below are constructed on the group $G=H_{3}$. By Corollary 1.8, we may assume that any infinite sequence $B\subset H_{3}$ considered below is infinite in all coordinates.
Example 6.1.
[…]
Consider the Følner sequence $\Phi=(\Phi_{N})_{N\in\mathbb{N}}$ with $\Phi_{N}=(2^{N}+[N])\times[N]\times[N^{2}]$. It is not hard to see that $\Phi$ is a left Følner sequence, but not a right one. Let $\Phi_{N}^{\prime}=(2^{N}+[N]^{\prime})\times[N]^{\prime}\times[N^{2}]^{\prime}$ and consider the set $A=\bigcup_{N\in\mathbb{N}}\Phi_{N}^{\prime}$. Clearly, $\overline{d}_{\Phi}(A)>0$. Suppose, for sake of contradiction, that there exist an infinite sequence $B=(b(n))_{n\in\mathbb{N}}\subset A$ and some $t\in H_{3}$ such that $\{b(i)b(j)\colon i<j\}\subset At^{-1}$. We denote $t^{-1}=(t_{1},t_{2},t_{3})$. We observe that $B\cap\Phi_{N}^{\prime}\neq\emptyset$ for infinitely many $N\in\mathbb{N}$. Let $b=b(i)=(b_{1},b_{2},b_{3})$, for some $i\in\mathbb{N}$, such that $b\in\Phi_{N}^{\prime}$ for some $N$ large (compared to the $t_{i}$’s). Then we can find some $j>i$ such that the element $c=b(j)=(c_{1},c_{2},c_{3})$ belongs in some $\Phi_{M}^{\prime}$ for some $M$ much larger than $N$, and then $bc\in At^{-1}$. It follows that $bc\in\Phi_{Q}^{\prime}t^{-1}$, for some $Q\in\mathbb{N}$, where
$\Phi_{Q}^{\prime}t^{-1}=(2^{Q}+[Q]^{\prime}+t_{1})\times([Q]^{\prime}+t_{2})\times([Q^{2}]^{\prime}+t_{3}+t_{2}(2^{Q}+[Q]^{\prime})).$
Then $(bc)_{1}={\rm O}(2^{Q})$ and on the other hand, $(bc)_{1}=b_{1}+c_{1}\in 2^{N}+2^{M}+[N+M]^{\prime}$, and hence we have that $(bc)_{1}={\rm O}(2^{M})$. It follows that $Q$ and $M$ are of the same order (and, in fact, we can show that they are equal). We now want to show that $t_{2}\neq 0$. By the fact that $bc\in\Phi_{Q}^{\prime}t^{-1}$, we have that
$bc\in(2\mathbb{Z}+1+t_{1})\times(2\mathbb{Z}+1+t_{2})\times(2\mathbb{Z}+1+t_{3}+t_{2}).$
On the other hand, multiplying $b$ and $c$, and using that $b_{i}$ and $c_{i}$ are odd for all $i$, gives that
$bc=(b_{1}+c_{1},b_{2}+c_{2},b_{3}+c_{3}+b_{1}c_{2})\in 2\mathbb{Z}\times 2\mathbb{Z}\times(2\mathbb{Z}+1).$
It follows that all the $t_{i}$’s are odd, and in particular, $t_{2}\neq 0$.
Now, since $cb\in\Phi_{Q}^{\prime}t^{-1}$ and $t_{2}\neq 0$, we have that
$(bc)_{3}\gg 2^{Q}\gg 2^{M}.$
On the hand, for $M$ sufficiently large, we have that
$(bc)_{3}=b_{3}+c_{3}+b_{1}c_{2}\leq N^{2}+M^{2}+(2^{N}+N)M\ll M^{2},$
which yields a contradiction.
Example 6.2.
[…]
Consider the same Følner sequence $\Phi$ as in Example 6.1. We observe that $\Phi_{N}\cap\Phi_{M}=\emptyset$ for any $N\neq M$, and in particular, the projections of any two such sets in the first coordinate are disjoint subsets of $\mathbb{Z}$. We define the set $A=\bigcup_{N\in\mathbb{N}}\Phi_{N}$ and clearly we have $\overline{d}_{\Phi}(A)=1$. Suppose, for sake of contradiction, that there exist an infinite sequence $B=(b(n))_{n\in\mathbb{N}}\subset H_{3}$ and some $t\in H_{3}$ such that $\{b(i)b(j)\colon i>j\}\subset t^{-1}A$. We denote $t=(t_{1},t_{2},t_{3})$ and $b(1)=b=(b_{1},b_{2},b_{3})$. We may assume without loss of generality that $b_{2}\neq 0$. We let $B^{\prime}=(b(n))_{n\geq 2}$. Moreover, we denote $b^{-1}=y=(y_{1},y_{2},y_{3})$ and then we have that $B^{\prime}\subset t^{-1}Ay$. It follows that $B^{\prime}\cap t^{-1}\Phi_{N}y\neq\emptyset$ for infinitely many $N\in\mathbb{N}$. Fix $c=b(i)$ for some $i>1$ such that $c\in B^{\prime}\cap t^{-1}\Phi_{M}y$ for some large $M\in\mathbb{N}$. Then $cb\in t^{-1}A$, which implies that $cb$ belongs in exactly one set of the form $t^{-1}\Phi_{N}$. We have that
$(cb)_{1}=c_{1}+b_{1}\in 2^{M}+[M]+t_{1}+y_{1}+b_{1}=2^{M}+[M]+t_{1},$
hence $cb\in t^{-1}\Phi_{M}$. Then we have that $(cb)_{3}\in[M^{2}]+t_{1}[M]+t_{3}$, which imples that $(cb)_{3}\ll M^{2}$, for $M$ sufficiently large. On the other hand, multiplying $c$ and $b$ gives that
$(cb)_{3}=c_{3}+b_{3}+c_{1}b_{2}\gg c_{1}\in 2^{M}+[M]+t_{1}+y_{1},$
which implies that $(cb)_{3}\gg 2^{M}=2^{M}$, for $M$ sufficiently large, where the implied constant is again absolute. This yields a contradiction.
Example 6.3.
[…]
To see why, suppose, for sake of contradiction, that there exists such a sequence $B$ and, as we did in Remark 1.13, consider the sequence $B^{\prime}=t^{-1}Bt$. Then we have that $B^{\prime}\hskip 2.84526pt\mathrlap{\mkern 0.9mu{{\cdot}}}{\triangleright}\hskip 2.13394ptB^{\prime}\subset t^{-1}A$, which cannot hold for this particular set $A$ as we saw in Example 6.2. This yields a contradiction.
Appendix A A result of Host and Kra for amenable groups
The purpose of this appendix is to prove Lemma 3.5. The proof follows the ideas in the proof of [HK09, Proof of Proposition 6.1], adapted in our setting. We state the following classical result (see for example [Rud91, Example 11.13 (a)]), which we will need in the proof of Lemma 3.5:
Lemma A.1.
Let $X$ be a compact metric space. Then the only linear multiplicative functionals on the algebra $C(X)$ are the point evaluations, i.e., $\textup{ev}_{x}(f)=f(x)$, for $x\in X$.
For convenience, we restate the lemma we want to prove:
Lemma 3.5.
[HK09, Proposition 6.1 for group actions] Let $G$ be an amenable group, let $(X,\mu,T)$ be an ergodic $G$-system, $(Z,m,R)$ be its Kronecker factor and $\rho:(X,\mu,T)\to(Z,m,R)$ be a factor map. If $a\in X$ is a transitive point, then there exists a point $z\in Z$ and a Følner sequence $\Psi$ such that
(A.1) $\lim_{N\to\infty}\frac{1}{|\Psi_{N}|}\sum_{g\in\Psi_{N}}f_{1}(T_{g}a)\cdot f_{2}(R_{g}z)=\int_{X}f_{1}\cdot(f_{2}\circ\rho)~{}\mathrm{d}\mu$
holds for any $f_{1}\in C(X)$ and $f_{2}\in C(Z)$.
We remark that the result still holds if we replace $(Z,m,R)$ by any factor of $(X,\mu,T)$ that is distal as a topological system.
Proof.
As in [HK09], we split the proof into two parts.
Construction of a common extension. Let $\mathcal{A}\subset\{f:X\to\mathbb{C}\colon f~{}\text{is measurable and bounded}\}$ be the closed (in norm) subalgebra that is spanned by $C(X)$ and $\{f\circ\rho:f\in C(Z)\}$. This is a unital commutative separable algebra, which contains the constants and is invariant under both complex conjugation and $T$. Consider the Gelfand spectrum of $\mathcal{A}$, which is defined as
$W=\{\chi:\mathcal{A}\to\mathbb{C}\colon\chi~{}\text{is linear and multiplicative}\}.$
Note that $W$ is compact and metrizable, since $\mathcal{A}$ is separable. By Gelfand’s theorem, there exists an isometric isomorphism $F:C(W)\to\mathcal{A}$, satisfying $F^{-1}(f)(\chi)=\chi(f)$ for all $f\in\mathcal{A}$ and all $\chi\in W$. Hence for all $\widetilde{f}\in C(W)$ and all $\chi\in W$, we have $\widetilde{f}(\chi)=\chi(F(\widetilde{f}))$. For $g\in G$ and $\chi\in W$, we define $S_{g}(\chi):A\to\mathbb{C}$, $S_{g}(\chi)(f)=\chi(f\circ T_{g})$. Then it is not too difficult to see that for each $g\in G$, $S_{g}:W\to W$ is a homeomorphism, and we also let $S=(S_{g})_{g\in G}$. Then for every $\chi\in W$, we have that
$\chi(F(\widetilde{f}\circ S_{g}))=\widetilde{f}(S_{g}(\chi))=S_{g}(\chi)(F(\widetilde{f}))=\chi(F(\widetilde{f})\circ T_{g})$
for $\widetilde{f}\in C(W)$ and any $g\in G$. In particular, for every $x\in X$, by considering the evaluation functional $\textup{ev}_{x}\in W$, it follows that
$F(\widetilde{f}\circ S_{g})(x)=\textup{ev}_{x}(F(\widetilde{f}\circ S_{g}))=\textup{ev}_{x}(F(\widetilde{f})\circ T_{g})=(F(\widetilde{f})\circ T_{g})(x)$
holds for any $\widetilde{f}\in C(W)$ and any $g\in G$. Thus, we have that
(A.2) $F(\widetilde{f}\circ S_{g})=F(\widetilde{f})\circ T_{g}$
for every $\widetilde{f}\in C(W)$ and for every $g\in G$.
Now, we consider the embedding $F^{-1}|_{C(X)}:C(X)\hookrightarrow C(W)$. Given $w\in W$, $\textup{ev}_{w}\circ F^{-1}|_{C(X)}$ is a linear multiplicative functional on $C(X)$, and by Lemma A.1, there exists a unique $x\in X$ such that $\textup{ev}_{w}\circ F^{-1}|_{C(X)}=\textup{ev}_{x}$. Thus, we can define $\pi_{X}:W\to X$ by $\pi_{X}(w)=x$ if and only if $\textup{ev}_{w}\circ F^{-1}|_{C(X)}=\textup{ev}_{x}$. The last equation is equivalent to that for any $f\in C(X)$ and $w\in W$,
$f\circ\pi_{X}(w)=\textup{ev}_{\pi_{X}(w)}(f)=\textup{ev}_{w}\circ F^{-1}|_{C(X)}(f)=F^{-1}(f)(w).$
Hence $\pi_{X}$ is the unique map from $W$ to $X$ satisfying
(A.3) $f\circ\pi_{X}=F^{-1}(f)$
for any $f\in C(X)$. We claim that $\pi_{X}$ is continuous and surjective.
To show continuity, we let $(w_{n})_{n\in\mathbb{N}}$ in $W$ such that $w_{n}\to w\in W$. Then for any $f\in C(W)$, we have $\textup{ev}_{w_{n}}(f)\to\textup{ev}_{w}(f)$. Hence, for any $f\in C(X)$, we have $\textup{ev}_{\pi_{X}(w_{n})}(f)\to\textup{ev}_{\pi_{X}(w)}(f)$, that is, $f(\pi_{X}(w_{n}))\to f(\pi_{X}(w))$. Since $X$ is compact, $\pi_{X}(w_{n})$ has a convergent subsequence, which by abuse of notation we denote by $\pi_{X}(w_{n})$. Suppose for sake of contradiction that $\pi_{X}(w_{n})\to y\neq\pi_{X}(w)$. Then by Urysohn’s lemma, we can find some $f\in C(X)$ and some disjoint open neighborhoods $U_{1}\ni\pi_{X}(w),U_{2}\ni y$, such that $f|_{U_{1}}=1$ and $f|_{U_{2}}=0$. It follows that $\pi_{X}(w_{n})\in U_{2}$ for large $n$, hence $f(\pi_{X}(w_{n}))=0$ for large $n$, while $f(\pi_{X}(w))=1$, but this contradicts the fact that $f(\pi_{X}(w_{n}))\to f(\pi_{X}(w))$. This shows that every convergent subsequence of $\pi_{X}(w_{n})$ converges to $\pi_{X}(w)$, and since $X$ is compact, it follows that $\pi_{X}(w_{n})\to\pi_{X}(w)$, showing the continuity of $\pi_{X}$.
To show that $\pi_{X}$ is surjective, let $x\in X$ and consider the linear multiplicative functional $\textup{ev}_{x}$ on $C(X)$. Since $F^{-1}|_{C(X)}$ is an embedding, it follows that $F^{-1}|_{C(X)}\circ\textup{ev}_{x}$ is a linear multiplicative functional on $C(W)$. Then, by Lemma A.1, there exists some $w\in W$ such that $F^{-1}|_{C(X)}\circ\textup{ev}_{x}=\textup{ev}_{w}$. Hence, $\pi_{X}(w)=x$, showing that $\pi_{X}$ is surjective.
Moreover, for any $g\in G$ and any $f\in C(X)$, let $\widetilde{f}=F^{-1}(f)\in C(W)$, and then $F(\widetilde{f})=f\in C(X)$ and by (A.2), $F(\widetilde{f}\circ S_{g})=f\circ T_{g}\in C(X)$. Then, using (A.2) and (A.3), we have that
$f\circ T_{g}\circ\pi_{X}=F(\widetilde{f}\circ S_{g})\circ\pi_{X}=\widetilde{f}\circ S_{g}=F(\widetilde{f})\circ\pi_{X}\circ S_{g}=f\circ\pi_{X}\circ S_{g}$
for any $g\in G$ and any $f\in C(X)$. It follows by Urysohn’s lemma that
(A.4) $T_{g}\circ\pi_{X}=\pi_{X}\circ S_{g}$
for any $g\in G$. Therefore, we have proved that $W$ is an extension of $X$ with $\pi_{X}$ being a continuous topological factor map.
Similarly, by considering the embedding $F^{-1}|_{C(Z)\circ\rho}:C(Z)\circ\rho\hookrightarrow C(W)$, there exists a unique surjective continuous map $\pi_{Z}:W\to Z$ such that
(A.5) $f\circ\pi_{Z}=F^{-1}(f\circ\rho)$
for any $f\in C(Z)$, and
(A.6) $R_{g}\circ\pi_{Z}=\pi_{Z}\circ S_{g}$
for any $g\in G$. Hence, $W$ is also an extension of $Z$ with $\pi_{Z}$ being a continuous topological factor map.
Now we will find a measure on $W$, with which $W$ will become a measurable extension of $X$ and $Z$. Since $f\mapsto\int f~{}\mathrm{d}\mu$ is a positive linear functional on $\mathcal{A}$, there exists a unique probability measure $\nu$ on $W$ such that
$\int_{X}f~{}\mathrm{d}\mu=\int_{W}F^{-1}(f)~{}\mathrm{d}\nu$
for any $f\in\mathcal{A}$. By (A.2), we have that $\nu$ is $S$-invariant, by (A.3), we have $\pi_{X}\nu=\mu$ and by (A.5), we have $\pi_{Z}\nu=m$. Consequently, $\pi_{X}$ and $\pi_{Z}$ are factor maps.
The last thing in this first step is to show that $\pi_{X}$ is actually a measurable isomorphism between $W$ and $X$ and thus, that the measure $\nu$ is ergodic. First, we want to extend (A.3) in $C(W)\simeq\mathcal{A}$. For $f\in\mathcal{A}$, it holds $\int_{W}|F^{-1}(f)|^{2}~{}\mathrm{d}\nu=\int_{X}|f|^{2}~{}\mathrm{d}\mu$, and $F^{-1}$ is an isometry from $\mathcal{A}$ (with the $L^{2}(X,\mu)$ norm) into $L^{2}(W,\nu)$. Combining the facts that $C(X)$ is dense in $\mathcal{A}$ (with respect to the $L^{2}(X,\mu)$ norm) and that (A.3) holds for all $f\in C(X)$, we obtain that (A.3) holds for all $f\in\mathcal{A}$, $\nu$-almost always. Then consider the map $H:L^{2}(X,\mu)\to L^{2}(W,\nu)$, such that $f\mapsto f\circ\pi_{X}$. Then $H(L^{2}(\mu))$ is closed in $L^{2}(W,\nu)$, since the map is an isometry, and notice that it contains $F^{-1}(\mathcal{A})=C(W)$. Thus, $H(L^{2}(X,\mu))=L^{2}(W,\nu)$, showing that $\pi_{X}$ is a measurable isomorphism, and consequently, that $(W,\nu,S)$ is ergodic. Finally, for any $f\in C(Z)$, using (A.3) and (A.5), we see that $f\circ\pi_{Z}=F^{-1}(f\circ\rho)=f\circ\rho\circ\pi_{X}$ holds $\nu$-almost always, and so, $\pi_{Z}=\rho\circ\pi_{X}$.
Construction of the Følner sequence. Since $(W,\nu,S)$ is ergodic, it follows that there exists $w_{1}\in\texttt{{gen}}(\nu,\Phi)$ for some Følner sequence $\Phi$.
Set $x_{1}=\pi_{X}(w_{1})$. Transitivity of $a$ implies that there exists a sequence $(h_{N})_{N\in\mathbb{N}}\subset G$ such that
(A.7) $\lim_{N\to\infty}\sup_{g\in\Phi_{N}}d_{X}(T_{g}x_{1},T_{gh_{N}}a)=0,$
where $d_{X}$ is the metric on the space $X$. Now set $z_{1}=\pi_{Z}(w_{1})$. Let $E(Z,R)$ be the Ellis semigroup of $(Z,R)$, that is, the closure of $R$ as an element of $Z^{Z}$, where this space is equipped with the pointwise convergence topology. Let $R_{0}\in\overline{(R_{h_{N}})}_{N\in\mathbb{N}}\subset E(Z,R)$. By Proposition 2.11 $R$ is a rotation, which implies that is a bijection from $Z$ to itself. In case that $(Z,m,R)$ is any distal system (and not necessarily the Kronecker factor), then we also have that $R$ is a bijection (see by [Aus88, Chapter 5]). Therefore, there exists $z_{0}\in Z$ such that $R_{0}(z_{0})=z_{1}$. Then there exists a subsequence of $(h_{N})_{N\in\mathbb{N}}$, which, by abuse of notation, we denote as $(h_{N})_{N\in\mathbb{N}}$, such that $\lim_{N\to\infty}R_{h_{N}}z_{0}=z_{1}$. Therefore, there exists a further subsequence, which once again we denote in the same way, for which it holds that
(A.8) $\lim_{N\to\infty}\sup_{g\in\Phi_{N}}d_{Z}(R_{g}z_{1},R_{gh_{N}}z_{0})=0,$
where $d_{Z}$ is the metric on the space $Z$.
Let $f_{1}\in C(X)$, $f_{2}\in C(Z)$. By (A.7) and (A.8), we have that
$\lim_{N\to\infty}\sup_{g\in\Phi_{N}}|f_{1}(T_{g}x_{1})-f_{1}(T_{gh_{N}}a)|=0\text{\quad and \quad}\lim_{N\to\infty}\sup_{g\in\Phi_{N}}|f_{2}(R_{g}z_{1})-f_{2}(R_{gh_{N}}z_{0})|=0.$
We define the Følner sequence $\Psi=(\Psi_{N})_{N\in\mathbb{N}}$, by $\Psi_{N}=\Phi_{N}h_{N}$ for any $N\in\mathbb{N}$. It is easy to check that since $\Phi$ is a left Følner sequence, then $\Psi$ is also a left Følner sequence. It follows from the above equations that
$\lim_{N\to\infty}\bigg{|}\frac{1}{|\Psi_{N}|}\sum_{g\in\Psi_{N}}f_{1}(T_{g}a)f_{2}(R_{g}z_{0})-\frac{1}{|\Phi_{N}|}\sum_{g\in\Phi_{N}}f_{1}(T_{g}x_{1})f_{2}(R_{g}z_{1})\bigg{|}=$
$=\lim_{N\to\infty}\bigg{|}\frac{1}{|\Phi_{N}|}\sum_{g\in\Phi_{N}}\big{(}f_{1}(T_{gh_{N}}a)f_{2}(R_{gh_{N}}z_{0})-f_{1}(T_{g}x_{1})f_{2}(R_{g}z_{1})\big{)}\bigg{|}$
$\leq\lim_{N\to\infty}\sup_{g\in\Phi_{N}}|f_{1}(T_{gh_{N}}a)f_{2}(R_{gh_{N}}z_{0})-f_{1}(T_{g}x_{1})f_{2}(R_{g}z_{1})|$
$\leq\lim_{N\to\infty}\sup_{g\in\Phi_{N}}\big{(}|f_{1}(T_{gh_{N}}a)f_{2}(R_{gh_{N}}z_{0})-f_{1}(T_{gh_{N}}a)f_{2}(R_{g}z_{1})|$
$\hskip 71.13188pt+|f_{1}(T_{gh_{N}}a)f_{2}(R_{g}z_{1})-f_{1}(T_{g}x_{1})f_{2}(R_{g}z_{1})|\big{)}$
$\leq\lim_{N\to\infty}\big{(}\|f_{1}\|_{\infty}\sup_{g\in\Phi_{N}}|f_{2}(R_{gh_{N}}z_{0})-f_{2}(R_{g}z_{1})|+\|f_{2}\|_{\infty}\sup_{g\in\Phi_{N}}|f_{1}(T_{gh_{N}}a)-f_{1}(T_{g}x_{1})|\big{)}$
(A.9) $=0.$
Moreover, recalling that $w_{1}\in\texttt{{gen}}(\nu,\Phi)$ and observing that $f_{1}\circ\pi_{X}\in C(W)$ and $f_{2}\circ\pi_{Z}\in C(W)$, we have that
$\lim_{N\to\infty}\frac{1}{|\Phi_{N}|}$ $\sum_{g\in\Phi_{N}}f_{1}(T_{g}x_{1})f_{2}(R_{g}z_{1})=\lim_{N\to\infty}\frac{1}{|\Phi_{N}|}\sum_{g\in\Phi_{N}}f_{1}(T_{g}(\pi_{X}(w_{1})))f_{2}(R_{g}(\pi_{Z}(w_{1})))$
[…]
$=\int_{W}(f_{1}\circ\pi_{X})(f_{2}\circ\pi_{Z})~{}\mathrm{d}\nu$
$=\int_{W}(f_{1}\circ\pi_{X})(f_{2}\circ\rho\circ\pi_{X})~{}\mathrm{d}\nu\hskip 63.44952pt\text{ (since }\pi_{Z}=\rho\circ\pi_{X})$
[…]
Combining (A) and (A) yields the desired result. The proof is complete. ∎